### v1

# What Language Models Know But Don’t Say: Non-Generative Prior Extraction for Generalization

Abstract
In domains like medicine and finance, large-scale labeled data is costly and often unavailable, leading to models trained on small datasets that struggle to generalize to real-world populations. Large language models contain extensive knowledge from years of research across these domains. We propose LoID (Logit-Informed Distributions), a deterministic method for extracting informative prior distributions for Bayesian logistic regression by directly accessing their token-level predictions. Rather than relying on generated text, we probe the model’s confidence in opposing semantic directions (positive vs. negative impact) through carefully constructed sentences. By measuring how consistently the LLM favors one direction across diverse phrasings, we extract the strength and reliability of the model’s belief about each feature’s influence. We evaluate LoID on ten real-world tabular datasets under synthetic out-of-distribution (OOD) settings characterized by covariate shift, where the training data represents only a subset of the population. We compare our approach against (1) standard uninformative priors, (2) AutoElicit, a recent method that prompts LLMs to generate expert-style priors via text completions, (3) LLMProcesses, a method that uses LLMs to generate numerical predictions through in-context learning by formatting tabular data as natural language prompts and (4) an oracle-style upper bound derived from fitting logistic regression on the full dataset. We assess performance using Area Under the Curve (AUC). Across datasets, LoID significantly improves performance over logistic regression trained on OOD data, recovering up to 59% of the performance gap relative to the oracle model. LoID outperforms AutoElicit and LLMProcessesc on 8 out of 10 datasets, while providing a reproducible and computationally efficient mechanism for integrating LLM knowledge into Bayesian inference.
What Language Models Know But Don’t Say:
Non-Generative Prior Extraction for Generalization
Sara Rezaeimanesh Michigan State University rezaeima@msu.edu          Mohammad M. Ghassemi Michigan State Unviersity ghassem3@msu.edu
1 Introduction
Machine learning models used in critical domains like medicine and finance often lack access to large, well-balanced, and comprehensively labeled datasets. Clinical datasets may cover only a single hospital or demographic, while credit-risk tables can be skewed toward historical lending patterns. For example, Purushotham et al. (2017) benchmarked several deep learning models on clinical prediction tasks—including mortality and length of stay—using the MIMIC-III intensive care electronic health record dataset from a single hospital site. As a consequence, models trained on small, non-representative samples frequently fail to generalize when confronted with real-world populations that differ from their training distribution Brigato and Iocchi (2020); Li et al. (2024); Althnian et al. (2021). Bayesian methods offer a principled way to encode prior knowledge and mitigate such brittleness, but manually eliciting high-quality priors at scale remains a longstanding challenge O’Hagan et al. (2006).
LLMs, trained on vast corpora of scientific literature and domain-specific data Brown et al. (2020), offer a promising new source of priors for downstream modeling. Recent efforts have explored LLMs both as zero-shot and few-shot predictors for tabular classification Hegselmann et al. (2023); Requeima et al. (2024), and as in-context learners capable of producing joint probabilistic predictions Shysheya et al. (2025). Other work has investigated the integration of LLMs with classical statistical methods: LLM-FE Abhyankar et al. (2025) uses LLMs to guide feature engineering, LLM-Lasso Zhang et al. (2025) for feature selection, and LLMCorr Zhong et al. (2024) for post-hoc model correction. AutoElicit Capstick et al. (2025) introduced a technique to query LLMs for Gaussian priors, but its reliance on generative sampling introduces run-to-run variability.
Parallel to this, out-of-distribution (OOD) generalization has gained increasing attention, as real-world deployment often involves data that differs from the development environment. Meta-learning approaches like TabPFN Hollmann et al. (2023) and Drift-Resilient TabPFN Helli et al. (2024) address OOD robustness through synthetic task generation and causal modeling. Others have proposed logit-based calibration strategies such as UE-NL Huang and Qiao (2023), or LLM-guided data augmentation methods like Summary Boosting Manikandan et al. (2023) and LLMOverTab Isomura et al. (2025) to cover underrepresented regions of the feature space. However, despite this momentum, few works have systematically explored how to extract priors from the internal structure of LLMs (e.g. hidden states, logits, etc.) for tabular tasks under distribution shift.
We address this gap with LoID (Logit-Informed Distributions), a deterministic framework that elicits informative priors for Bayesian logistic regression directly from an LLM’s token-level logits. Instead of relying on generated completions or examples, LoID probes the LLM with semantically paired sentences highlighting positive and negative feature-target relationships. The logit difference between the relational tokens quantifies directional belief, and the variance across paraphrased prompts yields an uncertainty estimate—together forming Normal priors over model coefficients. These priors integrate into Bayesian workflows, without any fine-tuning or in-context learning.
We evaluate LoID on ten publicly available tabular datasets that are widely used benchmarks and have appeared in numerous prior studies, ensuring comparability with existing methods. We compare LoID against uninformative Bayesian inference and LLMProcesses. To provide an upper bound, we additionally train and evaluate a logistic regression model on the entire dataset, offering insight into performance when full population data is available. Across extensive out-of-distribution splits, LoID outperforms existing approaches in predictive accuracy (AUC). By deterministically extracting domain knowledge encoded in large language models, LoID enables robust and reproducible inference.
In summary, our contributions include:
  • Logit-based deterministic prior elicitation. LoID is the first method to extract Bayesian priors directly from the token-level logits of large language models, eliminating the variability associated with text generation and prompt sampling.
  • Baseline comparison across ten OOD datasets. We evaluate LoID on diverse OOD tabular datasets from healthcare, finance, and social science. Our experiments demonstrate that LoID often outperforms existing approaches across most settings.
2 Related Work
2.1 LLMs for Tabular Data Modeling
The integration of LLMs into tabular data modeling has garnered attention, particularly in addressing challenges such as limited data availability, complex feature interactions, and the necessity for domain-specific feature engineering.
Several studies have explored the use of LLMs as standalone predictors for tabular data. Hegselmann et al. (2023) introduce TabLLM, a framework that applies large language models to zero-shot and few-shot classification of tabular data by converting rows into natural-language strings, enabling effective classification even with minimal labeled examples. Requeima et al. (2024) propose LLMProcesses, a framework that enables LLMs to produce probabilistic regression outputs by conditioning on both numerical data and natural language descriptions of prior knowledge. Shysheya et al. (2025) present JOLT, a method that leverages the in-context learning capabilities of LLMs to define joint probabilistic predictions over tabular data.
Beyond standalone applications, recent research has focused on leveraging LLMs to enhance classical statistical models. For instance, Abhyankar et al. (2025) introduced LLM-FE, a framework that combines evolutionary search with the reasoning capabilities of LLMs to iteratively generate and refine feature transformations for tabular data. Similarly, LLM-Lasso Zhang et al. (2025) is a domain-informed approach to feature selection and regularization by leveraging large language models to guide the selection process. Capstick et al. (2025) proposed AutoElicit, a method that interacts with LLMs to extract priors for Gaussian-based probabilistic models; however, this approach primarily focuses on Gaussian priors and lacks flexibility in handling other priors that may be more suitable for different data characteristics. In a different vein, Zhong et al. (2024) presented LLMCorr, a training-free framework where LLMs act as post-hoc correctors to propose corrections for the predictions of arbitrary machine learning models. Liu et al. (2023) investigated the reasoning capabilities of LLMs over tabular data and proposed combining symbolic and textual reasoning to improve performance on table-based question answering tasks.
2.2 Out-of-Distribution Generalization
OOD generalization is a fundamental challenge in real-world applications of tabular machine learning. Models trained on limited or biased datasets often fail to generalize when deployed on test data that differ from the training distribution. A variety of approaches have been proposed to address this issue through data augmentation, distributionally robust optimization, causal modeling, Bayesian inference, and most recently, the use of LLMs.
One line of work focuses on meta-learning for small-data generalization. Hollmann et al. (2023) introduced TabPFN, a foundation model trained on millions of synthetic tabular tasks to approximate Bayesian posterior predictive distributions in one forward pass. The model shows strong generalization across novel tasks and distributions. Building on this, Helli et al. (2024) proposed Drift-Resilient TabPFN, which introduces temporally shifting causal mechanisms during training to improve robustness under distribution shift.
Several recent studies have explored logit-based and uncertainty-aware methods for detecting or adapting to OOD inputs. Huang and Qiao (2023) proposed Uncertainty Estimation with Normalized Logits (UE-NL), a calibration technique that regularizes softmax confidence and improves both in-distribution accuracy and OOD detection.
In the LLM context, OOD robustness has been approached in several creative ways. Summary Boosting (Manikandan et al., 2023) uses GPT-generated feature summaries as weak learners to improve robustness in small or under-specified datasets. LLMOverTab (Isomura et al., 2025) performs data augmentation by prompting LLMs to generate realistic examples, improving classifier performance on underrepresented regions of the feature space. These strategies demonstrate that LLMs can inject external knowledge or expand the training distribution to mitigate generalization failure.
Our work, LoID, complements and extends this growing body of research by proposing the first deterministic logit-based method to elicit Bayesian priors from LLMs for structured tabular tasks. Our method consistently improves OOD generalization across ten tabular datasets, offering a lightweight and almost deterministic integration of LLM knowledge into classical statistical pipelines.
3 Method
We propose LoID (Logit-Informed Distribution), a deterministic framework for extracting priors for tabular dataset features from LLMs using token-level logits, rather than relying on generated samples. Our goal is to derive informative priors for Bayesian logistic regression that generalize robustly in low-data settings.
3.1 Problem Setting
Let $\mathcal{D}_{\text{train}}=\{(x_{i},y_{i})\}_{i=1}^{n}$ denote a training dataset, where each $x_{i}\in\mathbb{R}^{d}$ is a vector of $d$ features and $y_{i}\in\{0,1,\ldots,m\}$ is the corresponding label. In Bayesian logistic regression, we model the likelihood as
$p(y_{i}=t\mid x_{i},\boldsymbol{\beta})=\sigma(x_{i}^{\top}\boldsymbol{\beta}),$
where $\sigma(\cdot)$ denotes the sigmoid function and $\boldsymbol{\beta}\in\mathbb{R}^{d}$ represents the regression coefficients. Bayesian inference is performed by specifying a prior distribution $p(\boldsymbol{\beta})$ and computing the posterior $p(\boldsymbol{\beta}\mid\mathcal{D}_{\text{train}}).$ In data-scarce regimes, informative priors can substantially improve generalization. LoID constructs such priors by leveraging domain knowledge encoded in large language models, without access to the training data.
3.2 Querying LLM Beliefs via Logit Differences
For each feature $x_{j}$, we construct paired prompts that express opposing semantic claims about its effect on the target variable. Each prompt is carefully designed such that all relevant context appears before a sentiment-bearing token (e.g., “positive” or “negative”), ensuring that the token’s probability considers the entire context for the task, including the name of the feature and the prediction task (see examples in A.2.1). For example, given a prompt prefix like:
‘‘The impact of $f_{j}$ on $t$ is ’’
we use the probabilities from the LLM’s softmax over its vocabulary to extract the probability of the next token being “positive” or “negative”. We define:
$\mathcal{P}_{j}^{+}$ $=P(\texttt{``positive''}\mid\text{prefix})$
$\mathcal{P}_{j}^{-}$ $=P(\texttt{``negative''}\mid\text{prefix})$
where $f_{j}$ is a natural language description of feature $x_{j}$ (e.g., “age of the patient”) and $t$ describes the prediction task (e.g., “disease risk”).
We define the logit preference score for a single prompt $p_{j}$ as:
$\text{logit}(p_{j})=\log\left(\frac{p_{j}}{1-p_{j}}\right)$
Where
$p_{j}=\frac{P_{j}^{+}}{P_{j}^{+}+P_{j}^{-}}$
The logit preference score quantifies the extent to which the model prefers the token “positive” over “negative” in a given context. While this difference may also reflect other patterns learned during training, our hypothesis is that by constructing sentences that do not grammatically or syntactically favor either token, we reduce the influence of such confounding factors. As a result, the observed preference is driven primarily by the semantic relationship between the feature and the target concept. Finally, applying the logit transformation maps probabilities from the bounded interval (0, 1) to the unbounded real line, making the resulting scores directly appliable as coefficients in logistic regression models.
The mean prefrence score $\mu_{j}$ for feature $x_{j}$ is given by averaging this score across a set of paraphrased prompts T{T}:
$\mu_{j}=\frac{1}{|T|}\sum_{t\in T}\text{logit}_{t}(p_{j})$
This value reflects the LLM’s normalized preference for a positive versus negative association with the feature across diverse prompt formulations. Figure 1 demonstrates how LoID extracts informative priors from model logits.
Figure 1: LoID Pipeline: To extract informative priors, LoID uses multiple sentence templates for each feature and computes the probabilities of the tokens "positive" and "negative" as the next token. From these probabilities, we calculate the logit preference score, which quantifies the model’s preference for one token over the other. Using the logit preference scores from all sentences, we compute their mean and standard deviation. We then apply a transformation to the standard deviation to reduce overconfidence, allowing the data to have more influence on the resulting distribution.
3.3 Constructing Priors from Logits
We model each coefficient $\beta_{j}$ with a Gaussian prior: $\beta_{j}\sim\mathcal{N}(\mu_{j},\sigma_{j}^{2})$. The mean $\mu_{j}$ is the average logit preference score across paraphrases, which reflects the direction and strength of influence the LLM assigns to feature $x_{j}$. To capture uncertainty in the LLM’s belief, we define the standard deviation $\sigma_{j}$ as a scaled function of variability in the individual logit preference scores:
$\sigma_{j}^{2}=\alpha+\gamma\cdot\text{std}(logit_{1}(p_{j}),\ldots,logit_{|T|}(p_{j}))$
Here, $logit_{t}(p_{j})$ denotes the logit preference score for prompt $t$ and feature $j$. $\alpha>0$ and $\gamma>0$ are fixed scaling hyperparameters that can be tuned to optimize LoID’s performance on a certain task. We perform a comprehensive hyperparameter tuning in A.2.4 to find the hypermarameters that achieve the highest cumulative AUC. This design assigns higher variance to features for which the LLM expresses inconsistent preferences across paraphrases, thereby reflecting greater uncertainty about the feature’s influence in the model.
3.4 Bayesian Inference with LoID Priors
With the LoID priors in place, posterior inference is performed using the No-U-Turn Sampler (NUTS), a variant of Hamiltonian Monte Carlo (HMC). HMC augments the parameter space with auxiliary momentum variables and explores the posterior by simulating Hamiltonian dynamics, enabling efficient, gradient-informed proposals that avoid the random-walk behavior of traditional MCMC methods.
Specifically, HMC introduces a momentum variable $\mathbf{r}$ for the regression coefficients $\boldsymbol{\beta}$ and defines the Hamiltonian
$\mathcal{H}(\boldsymbol{\beta},\mathbf{r})=-\log p(\boldsymbol{\beta}\mid\mathcal{D}_{\text{train}})+\frac{1}{2}\mathbf{r}^{\top}\mathbf{r}.$
Proposals are generated by numerically simulating the resulting Hamiltonian dynamics using a leapfrog integrator, which leverages gradients of the log-posterior to propose distant yet highly probable states.
NUTS eliminates the need to manually tune the trajectory length by adaptively extending the simulation until a U-turn condition is detected, ensuring efficient exploration while maintaining detailed balance. This results in faster mixing and improved effective sample sizes, particularly in high-dimensional parameter spaces.
We run multiple NUTS chains with warm-up steps for step-size and mass-matrix adaptation, and discard burn-in samples before collecting posterior draws. All reported results are computed using posterior expectations estimated from these samples.
3.5 Summary
LoID transforms the implicit domain knowledge encoded in large language models into explicit, feature-level priors suitable for probabilistic modeling. Unlike prior approaches that rely on generated samples, in-context learning, or fine-tuning, LoID achieves this deterministically by directly probing token-level logits. This design enables reproducibility and principled integration into Bayesian models.
3.6 Evaluation
To assess the quality of the resulting posteriors, we compute the AUC score on the entire dataset using models trained solely on the OOD training split. We then compare this AUC to that of a logistic regression trained on the entire dataset. This comparison quantifies how much of the predictive performance gap, caused by training on an OOD distribution, is recovered by our method.
4 Experimental Setup
4.1 Datasets
We evaluate and compare LoID’s performance on ten publicly available tabular datasets for binary and multi-class classification: The Heart Disease (Cleveland) dataset Janosi et al. (1989) predicts the presence of heart disease based on clinical and demographic attributes such as cholesterol level, and blood pressure. The Indian Liver Patient Dataset (ILPD) Ramana and Venkateswarlu (2022) classifies whether a patient has liver disease using demographic information and blood test measurements. The Blood Transfusion Service Center dataset Yeh (2008a) predicts whether a donor will donate blood again based on historical donation behavior. The Room Occupancy Detection dataset Candanedo (2016b) determines whether a room is occupied using time-series environmental sensor readings.
The Diabetes dataset Islam et al. (2020) predicts the likelihood of diabetes onset based on self-reported symptoms. The Pima Indians Diabetes dataset Smith et al. (1988) predicts the presence of diabetes using diagnostic measurements such as glucose concentration and BMI. The Give Me Some Credit (Good Credit Score) dataset Fusion and Cukierski (2011) assesses an individual’s creditworthiness based on income, debt ratio, and delinquency history. The Bank Marketing dataset Moro et al. (2014) predicts whether a client will subscribe to a term deposit based on demographic attributes and marketing interactions.
The Adult Income (Census Income) dataset Becker and Kohavi (1996b) predicts whether an individual’s annual income exceeds \$50K using demographic and educational features. Finally, the Jungle Chess (Jungle) dataset Shapiro (1983) predicts game outcomes based on structured representations of board states, providing a symbolic decision-making task.
For each dataset, we construct synthetic OOD splits as described in Section A.4.2. To verify that the resulting training sets do not reflect the true data distribution, a logitic regression model is trained separately on the out-of-distribution and in-distribution training splits and evaluated on the entire dataset. As shown in Table 1, the model trained on the OOD train set consistently underperforms compared to the model trained on the full dataset. This confirms that generalization is impaired and the OOD train set does not present the full distribution of the dataset as well as the in distribution train set.
Across all datasets, we apply the same preprocessing pipeline: categorical features are one-hot encoded, and missing values are imputed with zeros. For datasets with an excessive number of features—namely Give Me Some Credit, and Adult Income—we perform feature selection using GPT-4o by providing feature descriptions and prompting it to return a list of the 15 most relevant features for the prediction task. The same reduced feature set is used consistently across all experimental conditions to ensure comparability.
Dataset AUC (IID) AUC (OOD)
Heart Disease 0.93 0.86
Liver Disease 0.77 0.70
Blood Donation 0.75 0.58
Room Occupancy 1.00 0.98
Diabetes 0.84 0.81
Income 0.91 0.78
Credit 0.79 0.71
Jungle 0.81 0.71
Pima Diabetes 0.84 0.81
Bank Marketing 0.89 0.51
Table 1: AUC comparison of logistic regression on IID vs. OOD test splits. Lower performance on OOD indicates distribution shift.
We designate the Heart Disease, Blood Donation, and Diabetes datasets as development datasets for evaluating prompt structures and hyperparameter tuning. The remaining five datasets are reserved for testing the generalization performance of the final LoID method. See (A.2.4) for details on the hyperparameter search process and value ranges explored.
Dataset OOD-LR AutoElicit (baseline) LLMProc (baseline) LoID (our method) Cap (upper bound) Gap Closed%
Heart Disease 0.87 0.92 0.86 0.90 0.93 +50
Liver Disease 0.70 0.36 0.60 0.69 0.77 -19
Blood Donation 0.58 0.71 0.63 0.67 0.75 +52
Room Occupancy 0.98 0.97 0.80 1.00 1.00 +78
Diabetes 0.81 0.83 0.70 0.84 0.84 +99
Good Credit Score 0.71 0.71 0.30 0.76 0.79 +60
Income 0.78 0.85 0.51 0.87 0.91 +57
Pima Diabetes 0.82 0.83 0.74 0.84 0.84 +88
Bank Marketing 0.51 0.51 0.51 0.72 0.89 +56
Jungle Chess 0.71 0.78 0.62 0.79 0.81 +70
Table 2: Comparison of AUC scores across datasets. “OOD-LR” reports the performance of logistic regression trained on an OOD training set. LoID results are obtained using priors from Gemma-2-27B, with $\alpha=0.2$, $\gamma=2.0$ and $N_{sent}=10$. Among Uniformative, AutoElicit, LLMProcesses, and LoID, the highest AUC for each dataset is shown in bold. The Cap condition serves as an oracle-style upper bound. All values are rounded to two decimal places. In some cases, LoID and Cap appear to have similar performance; however, the gap is not fully closed. This is due to rounding. For example, on Diabetes, LoID achieves 0.839 while Cap achieves 0.842.
4.2 LLM-Based Prior Extraction
We use the Gemma-2-27B, Gemma-3-27B, llama-3-8B, and GPT-2 large language models to extract logit-based priors as described in Section 3.2. We use these models in zero-shot mode without any fine-tuning or in-context examples. In our experiments, the queried lexical items are represented as single tokens across models, enabling direct logit-based comparisons.
Each dataset is processed independently. For each feature, we issue 10 prompts to extract token-level log-probabilities for the positive and negative completions. The number of prompts was selected via hyperparameter tuning (Section A.2.4). The resulting feature-wise prior means and variances are then used to define informative Gaussian priors in a Bayesian logistic regression model. Posterior inference is performed using PyMC with a Laplace approximation.
We compare LoID to the following baselines:
  • OOD-LR: Logistic regression model trained on OOD train set.
  • AutoElicit: LLM-generated expert-style priors via sampled completions.
  • LLMProc (LLMProcesses): In-context learning approach that generates numerical predictions by formatting tabular data as natural language prompts. For fair comparison, we use the OOD training examples as few shot examples for the LLM.
  • Cap: An oracle-style upper bound achieved by training a logistic regression model on the entire dataset, simulating access to the full population.
We evaluate each method using AUC score obtained from training on the ood train set and testing on the entire dataset.
5 Results
Table 2 summarizes the performance of each method across ten tabular datasets. Based on our experiments in Section A.2.2, Gemma-2-27B achieves the highest cumulative AUC among the evaluated LLMs. Therefore, Table 2 reports results obtained using this model.
5.1 AUC Comparison
LoID outperforms logistic regression trained on the OOD training set for all datasets except Liver Disease. Notably, on the Diabetes and Pima Diabetes datasets, LoID almost completely closes the performance gap induced by OOD training, indicating that it can successfully infer the underlying data distribution. On average, LoID closes 59% of the performance gap introduced by the OOD training setting.
LLMProcesses often underperforms even compared to uninformative priors. Since this approach does not explicitly produce priors, it is difficult to determine the exact cause of its poor performance. One possible explanation is its complete reliance on LLMs without leveraging traditional machine learning models, which are often more robust at pattern extraction in structured data.
LoID also achieves superior performance on 8 out of 10 datasets, achieving an average improvement of 99.7% in gap closed when it outperforms AutoElicit. To further understand these performance differences, we analyze the elicited prior distributions produced by both methods across selected winning and losing datasets. Section A.5.1 presents this comparison.
Finding 1: Non-Monotonic Relationship Between Prior Uncertainty and Performance
  • Optimal range ($\sigma\approx 0.4$$0.5$): Achieves best performance on 7 datasets: Liver, Income, Bank, Creditg, Diabetes, Pima, Occupancy
  • Too weak ($\sigma\gtrsim 0.8$): Underperforms AutoElicit: Blood, Heart
This shows that moderately weak priors ($\sigma\approx 0.4$$0.5$) provide sufficient regularization while allowing the Bayesian posterior to correct systematic biases in the elicited means through the data likelihood. Excessively weak priors ($\sigma\gtrsim 0.8$) fail to leverage LLM knowledge effectively, approaching uninformed priors.
LoID excels when Prior uncertainty is moderate, datasets have sufficient samples to correct sign biases, and AutoElicit’s tight priors impose incorrect strong beliefs.
Finding 3: Superior Mean Estimation Despite Directional Bias
While LoID exhibits systematic positive bias in sign coverage (0–6% negative priors vs. AutoElicit’s 20–50%), the absolute quality of the mean estimates appears superior. Evidence from performance gaps suggests:
  1. 1.
    Magnitude accuracy over sign accuracy: On datasets where LoID wins, the improvement over AutoElicit indicates that the LLM-elicited means—despite sign bias—capture feature importance better than AutoElicit’s conversational approach.
  2. 2.
    Comparison to uninformed priors: LoID achieves  145% better gap closed vs. uninformed priors (see full comparison in A.6.1), demonstrating that the mean estimates contain genuine signal beyond just the regularization effect of uncertainty.
  3. 3.
    Interaction with uncertainty: The combination of (a) approximately correct magnitudes in means and (b) moderate uncertainty allows the Bayesian posterior to correct sign errors while preserving magnitude information. In contrast, AutoElicit’s tight uncertainties amplify errors in both magnitude and direction.
5.2 Computational Efficiency.
[…]
6 Discussion
The success of our LoID method challenges the conventional assumption that more accurate prior means necessarily lead to better performance. AutoElicit provides strong, confident priors ($\sigma\approx 0.05$$0.12$) with $\sim$40–50% correct negative signs. However, the incorrect confident priors harm inference by overpowering the data likelihood.
On the other hand, LoID provides weaker and uncertain priors ($\sigma\approx 0.4$$0.5$). This allows the weaker means to be overridden by data, providing soft regularization: the priors gently guide inference toward reasonable parameter spaces without rigidly enforcing incorrect beliefs. The result is that even when LoID’s directional beliefs are wrong, the posterior can recover because the prior is not insisting on its incorrectness. However, the benefits of this approach diminish when priors become too weak ($\sigma>0.8$).
7 Limitations and Future Work
While our results demonstrate the effectiveness of LoID, several limitations should be noted, which also point toward directions for future work.
Semantic granularity.
Our current approach probes binary directional relationships (positive vs. negative). Extending this framework to multi-token or scalar outputs could capture more nuanced associations, such as non-monotonic effects.
Hyperparameter selection:
Although we use fixed hyperparameters ($\alpha=0.2$, $\gamma=2.0$) across all test datasets, optimal values vary by task. Developing automated calibration methods or meta-learning approaches for selecting these parameters would improve out-of-the-box applicability.
LLM access requirements:
LoID requires access to token-level logits, limiting compatibility with closed-source APIs.
8 Conclusion
In this work, we introduce LoID, a deterministic method for extracting priors from LLMs. We show that priors elicited with LoID substantially mitigate the performance degradation caused by training on small, out-of-distribution datasets, recovering over 59% of the lost performance.
Our comparative analysis with AutoElicit demonstrates that the effectiveness of the LoID approach stems from honest uncertainty quantification, and correct mean magnitude estimation for logistic regression coefficients. By maintaining moderate prior uncertainty ($\sigma\approx 0.4$$0.5$), LoID provides soft regularization while allowing the Bayesian posterior to correct systematic biases through the data likelihood.
As a result, LoID outperforms AutoElicit on 80% of benchmark datasets, achieving an average improvement of 99.7% in gap closed when it wins. Against LLMProcesses, LoID achieves superior performance on all 10 datasets, with particularly dramatic gains on datasets where LLMProcesses’s end-to-end LLM approach fails to leverage structured data patterns.
Appendix A Appendix
A.1 Reproducibility
Code Availability
To ensure reproducibility, we will release our full codebase, including detailed instructions for running the experiments and reproducing all reported results. We also provide the the posterior files generated by our experiments, as well as the scripts used to generate all figures and tables in the paper.
Data Availability
All datasets used in this study are publicly available through open-access repositories. Further details and references are provided in Section A.4. All preprocessing steps required to obtain the final train/test splits used in our experiments, along with the preprocessed data, are available in our code release.
Results Variability
To assess the stability and determinism of our approach, all experiments were repeated five times. The resulting 95% confidence intervals for AUC were effectively zero (0.000), indicating that our method yields highly consistent results across runs.
A.2 Hyperparameter Optimization
We used three datasets—Blood Donation, Heart Disease, and Diabetes—as development datasets to tune hyperparameters based on AUC score.
A.2.1 Number of Prompt Templates
We tested our method with 1–20 prompt templates analyzed how the number of templates affects LoID’s performance. The following are some examples of our prompt templates:
  1. 1.
    The impact of on is …
  2. 2.
    The relationship between and is …
  3. 3.
    The role of in is …
  4. 4.
    When considering , the effect on is …
  5. 5.
    The correlation between and is …
As shown in Figure 2, performance gains plateau after approximately 10 sentences. Around this point, performance also becomes more stable, while incorporating additional sentences incurs increased computational cost. Accordingly, we use 10 sentences in all experiments.
Figure 2: Hyperparameter ablation results demonstrating that prior uncertainty (beta) is the critical factor for performance, with F-statistics of 346.7 (blood), 47.3 (heart), and 583.2 (diabetes), while variance sensitivity (alpha) and number of sentence formulations (num_sentences) show no significant effect.
A.2.2 Model Selection
We compared the performances of LLMs–GPT-2, LLaMA 3.1 8B, Gemma-3-27B, and Gemma-2-27B in terms of AUC. The best averge performance was achieved with Gemma-2-27B, as shown in Table 3.
dataset blood diabetes heart average
model
gemma-3 0.65 0.84 0.89 0.79
gemma-2 0.67 0.84 0.90 0.80
llama-3-8b 0.56 0.80 0.85 0.73
gpt-2 0.53 0.82 0.84 0.71
Table 3: Model comparison across datasets using $\alpha=0.2$, $\gamma=2.0$, and 10 sentence formulations.
A.2.3 Variance Estimation Formula
The prior standard deviation $\sigma_{j}$ for feature $j$ quantifies the uncertainty in the elicited prior mean. We explored two methods for computing $\sigma_{j}$ from LLM token probabilities, each capturing different aspects of elicitation uncertainty.
Method 1: Variance of Logit Differences
This method uses the variance across multiple sentence formulations as a measure of prior uncertainty:
$\sigma_{j}^{2}=\alpha+\gamma\cdot\text{std}(logit_{1}(p_{j}),\ldots,logit_{|T|}(p_{j}))$
where
$\text{logit}(p_{j})=\log\left(\frac{p_{j}}{1-p_{j}}\right)$
and
$p_{j}=\frac{P_{j}^{+}}{P_{j}^{+}+P_{j}^{-}}$
Rationale: Higher variance across sentence formulations indicates inconsistent LLM responses, suggesting greater epistemic uncertainty. The hyperparameters $\alpha$ and $\gamma$ control the base uncertainty level and sensitivity to variance, respectively.
Method 2: Entropy-Based Uncertainty
This method uses Shannon entropy to measure the uncertainty in the LLM’s token probability distribution:
$\sigma_{j}=\beta\cdot\frac{1}{N_{\text{sent}}}\sum_{i=1}^{N_{\text{sent}}}H(p_{\text{pos}}^{(i)},p_{\text{neg}}^{(i)})$
Rationale: Higher entropy (closer to $\log 2\approx 0.693$) indicates the LLM is uncertain between “positive” and “negative” tokens, reflecting ambiguity about the feature’s effect direction. Lower entropy indicates confident predictions. The scaling factor $\beta$ maps entropy to prior standard deviation.
Results
Table 4 compares the variance estimation methods described above. While the overall differences in performance are small, the variance of logit differences consistently performs slightly better across all three datasets. Accordingly, we adopt this method for LoID.
Dataset variance Entropy $\Delta$
Blood 0.67 0.66 +0.01
Heart 0.90 0.89 +0.01
Diabetes 0.84 0.83 +0.01
Average 0.80 0.79 +0.01
Table 4: Comparison of variance of logit differences (used by LoID) vs. entropy-based uncertainty for uncertainty estimation.
A.2.4 Tuning $\alpha$, $\gamma$, and $N_{\text{sent}}$
As detailed in Section 3.3, we use the following formula to calculate standard deviation:
$\sigma_{j}^{2}=\alpha+\gamma\cdot\text{std}(logit_{1}(p_{j}),\ldots,logit_{|T|}(p_{j}))$
We searched over $\alpha\in\{0.1,0.2,0.3,0.5\}$, $\gamma\in\{1,2,3,4\}$, and $N_{\text{sent}}\in\{5,10,15,20\}$. Table 5 reports the hyperparameter settings that achieve the best performance for each evaluation dataset. As shown in the table, no single hyperparameter configuration generalizes across all datasets; instead, each dataset requires task-specific tuning to achieve optimal performance. Among the evaluated $\alpha$$\gamma$ combinations, $\alpha=0.2$, $\gamma=2.0$, and $N_{\text{sent}}=10$ yield the best average performance. For simplicity, we adopt these values for all remaining experiments.
Dataset Best AUC $\alpha$ $\beta$ $N_{\text{sent}}$
Blood 0.6704 0.2 4.0 10
Heart 0.8987 0.1 2.0 20
Diabetes 0.8392 0.5 1.0 5
Overall (avg) 0.8027 0.2 2.0 10
Table 5: Best hyperparameter configurations identified for each dataset in the ablation study.
Figure 2 illustrates the relationship between AUC and the hyperparameters $\alpha$, $\gamma$, and the number of sentences. Among these, $\gamma$ has the most pronounced effect on performance, with its influence on AUC appearing approximately linear. For the Blood Donation dataset, AUC increases as $\gamma$ increases, whereas for the Heart Disease and Diabetes datasets, AUC decreases with larger values of $\gamma$. These trends indicate that prior certainty plays a pivotal role in model performance. In contrast, we observe no consistent relationship between $\alpha$, the number of sentences, and AUC across datasets, suggesting that these factors have a limited impact on performance.
Consistent with the previous findings, analysis of variance (ANOVA) reveals that $\gamma$ is the only statistically significant predictor of out-of-distribution performance across all datasets (Table 6). The baseline uncertainty parameter $\gamma$ exhibits strong effects ($p<0.0001$). In contrast, both $\alpha$ and $N_{\text{sent}}$ show no significant impact ($p>0.05$) on any dataset.
Dataset Parameter F-statistic p-value
Blood $\alpha$ 0.231 0.8743
$\gamma$ 346.743 <0.0001<0.0001
$N_{\text{sent}}$ 0.436 0.7280
Heart $\alpha$ 0.360 0.7820
$\gamma$ 47.269 <0.0001<0.0001
$N_{\text{sent}}$ 1.174 0.3285
Diabetes $\alpha$ 1.162 0.3331
$\gamma$ 583.151 <0.0001<0.0001
$N_{\text{sent}}$ 0.083 0.9691
Table 6: ANOVA results testing hyperparameter effects on out-of-distribution AUC. Only $\gamma$ (baseline prior uncertainty) shows significant effects across all datasets.
A.3 Additional Experiments
A.3.1 Feature Synonyms
We evaluated the impact of feature synonyms on model performance to determine whether changing feature names could introduce bias based on word selection. To do this, we created three sets of synonyms for the features in our evaluation datasets. Table 7 presents examples of these features and their corresponding synonyms.
Dataset Original Feature Synonym Set 1 Synonym Set 2
Blood recency Time since last donation Days elapsed since previous donation
frequency Number of donations made Donation count
Heart Age Patient age Years of life
Cholesterol Serum cholesterol level Blood cholesterol concentration
MaxHR Peak heart rate Highest heart rate reached
ExerciseAngina Angina during exercise Exercise-provoked chest pain
Diabetes Glucose Blood glucose level Plasma glucose concentration
BMI Body mass index Weight-to-height ratio
DiabetesPedigree Family diabetes history score Genetic diabetes risk
Table 7: Representative examples of feature name variations tested in synonym ablation study (Synonym Set 0 = original feature names).
Results in Table 8 reveal minimal performance variation across three distinct feature naming schemes, with AUC ranges of 0.007, 0.002, and 0.001 for blood, heart, and diabetes datasets respectively. These negligible differences demonstrate that our elicitation method is robust to superficial linguistic variations in feature descriptions. This stability is significant because it indicates the LLM is capturing genuine semantic understanding of the features’ predictive relationships rather than relying on brittle pattern matching or memorized associations with specific phrasings.
Dataset Syn0 Syn1 Syn2 Range
Bank 0.692 0.692 0.696 0.004
Blood 0.657 0.654 0.655 0.004
Credit Score 0.755 0.755 0.755 0.000
Diabetes 0.836 0.835 0.835 0.000
Heart 0.887 0.887 0.887 0.000
Income 0.844 0.842 0.844 0.002
Liver 0.667 0.666 0.667 0.001
Occupancy 0.988 0.988 0.988 0.001
Jungle 0.79 0.778 0.778 0.012
Pima 0.836 0.836 0.837 0.000
Mean Range 0.0013
Table 8: Synonym Ablation Study: Robustness to feature name variations across all datasets. Three synonym sets are tested.
A.3.2 Feature Description
We analyzed the impact of adding further description to the features and the target could help the LLM arrive at better estimates of the prior distributions.
Table 9 reveals minimal and inconsistent impact from including feature names alongside descriptions in prompts, with a mean absolute AUC difference of only 0.0084 (0.84%) across all 10 datasets. Performance changes are bidirectional, with 5 datasets showing marginal improvements (range: +0.002 to +0.038) and 5 showing marginal declines (range: -0.003 to -0.016) when feature descriptions are included. This 50/50 split and small standard deviation (0.0135) indicate no systematic benefit or detriment to either prompt format, suggesting the LLM extracts equivalent semantic information regardless of whether features are referenced by name-description pairs or names alone. The method’s robustness to this formatting choice demonstrates that its effectiveness stems from semantic understanding rather than sensitivity to linguistic structure.
Dataset w/o Desc w/ Desc $\Delta$
Bank 0.692 0.673 -0.019
Blood 0.652 0.650 -0.002
Credit Score 0.756 0.756 +0.000
Diabetes 0.836 0.836 +0.000
Heart 0.886 0.893 +0.007
Income 0.845 0.853 +0.008
Liver 0.668 0.663 -0.005
Occupancy 0.989 0.986 -0.002
Jungle 0.778 0.777 -0.000
Pima 0.837 0.835 -0.002
Mean $|\Delta|$ 0.0050
Table 9: Feature Description Ablation Study: Impact of including feature descriptions in prompts across all datasets. Comparing name-only prompts vs. name+description format.
A.4 Datasets and OOD Splits
A.4.1 Datasets
We evaluate our prior elicitation method on ten binary, and multi-class classification datasets:
  1. 1.
    Bank Marketing Moro et al. (2014): Predicts whether a client will subscribe to a term deposit based on direct marketing campaign data from a Portuguese banking institution. Features include client demographics (age, job, marital status, education), campaign information (contact type, number of contacts, previous campaign outcome), and economic indicators (employment variation rate, consumer price index). The binary classification task identifies clients likely to subscribe to term deposits.
  2. 2.
    Blood Transfusion Yeh (2008b): Predicts whether a blood donor will donate within a given time window using data from the Blood Transfusion Service Center in Hsin-Chu City, Taiwan. Features are based on the RFMTC marketing model: Recency (months since last donation), Frequency (total number of donations), Monetary (total blood donated in c.c.), and Time (months since first donation). The binary classification task identifies donors likely to donate in March 2007.
  3. 3.
    Give Me Some Credit Kaggle (2011): Predicts the probability of financial distress in the next two years using credit and demographic data from Kaggle. Features include checking account status, credit duration, credit history, credit amount, savings account balance, employment status, and age. The binary classification task evaluates financial risk for 1,000 borrowers.
  4. 4.
    Diabetes Islam et al. (2020): Predicts the likelihood of diabetes based on self-reported symptoms and demographic information. Features include age, gender, polyuria, polydipsia, sudden weight loss, weakness, polyphagia, genital thrush, visual blurring, itching, irritability, delayed healing, partial paresis, muscle stiffness, alopecia, and obesity. The dataset contains 520 instances collected via questionnaires and is formulated as a binary classification task.
  5. 5.
    Heart Disease Janosi et al. (1988): Predicts heart disease using clinical and demographic attributes such as age, sex, resting blood pressure, cholesterol level, fasting blood sugar, maximum heart rate, exercise-induced angina, ST depression (oldpeak), chest pain type, resting ECG results, and ST slope. The binary classification task identifies individuals with high likelihood of heart disease.
  6. 6.
    Adult Income Becker and Kohavi (1996a): Predicts whether an individual’s annual income exceeds \$50K based on 1994 U.S. Census data. Features include age, education level, marital status, occupation, relationship status, workclass, sex, native country, capital gain/loss, and hours worked per week. The binary classification task uses demographic and employment data from 48,842 records.
  7. 7.
    Chess Endgame Shapiro (1983): Predicts game outcomes in King-Rook vs. King-Pawn chess endgame positions. Features encode piece positions using strength rankings, file coordinates, and rank coordinates for white and black pieces. The binary classification task determines winning positions in this chess endgame with 44,820 board configurations.
  8. 8.
    Indian Liver Patient Ramana and Venkateswarlu (2012): Predicts whether a patient has liver disease using clinical measurements collected from northeast Andhra Pradesh, India. Features include age, gender, total bilirubin, direct bilirubin, alkaline phosphatase, alanine aminotransferase (SGPT), aspartate aminotransferase (SGOT), total proteins, albumin, and albumin/globulin ratio. The binary classification task distinguishes liver patients from healthy individuals using data from 583 records (416 liver patients, 167 non-liver patients).
  9. 9.
    Occupancy Detection Candanedo (2016a): Predicts room occupancy status from environmental sensor measurements in an office setting. Features include temperature (Celsius), relative humidity (%), light (Lux), CO2 concentration (ppm), humidity ratio (kg water-vapor/kg-air), and timestamp information. The binary classification task detects occupancy using measurements recorded at one-minute intervals.
  10. 10.
    Pima Indians Diabetes Smith et al. (1988): Predicts the onset of diabetes within five years in females of Pima Indian heritage over age 21. Features include number of pregnancies, glucose level, blood pressure, skin thickness, insulin level, BMI, diabetes pedigree function, and age. The binary classification task uses medical measurements from 768 women from a population near Phoenix, Arizona with high diabetes incidence rates.
Dataset Method Mean Mag. Std (Uncertainty) Std Ratio Performance
[…]
AutoElicit 0.051 0.058 $1\times$
[…]
AutoElicit 0.015 0.124 $1\times$
[…]
AutoElicit 0.017 1.379 $1\times$
[…]
AutoElicit 0.008 0.047 $1\times$
[…]
AutoElicit 0.015 0.087 $1\times$
[…]
AutoElicit 0.070 0.044 $1\times$
[…]
AutoElicit 0.085 0.059 $1\times$
[…]
AutoElicit 0.007 0.122 $1\times$
[…]
AutoElicit 0.025 0.114 $1\times$
[…]
AutoElicit 0.041 0.024 $1\times$
Table 10: Comparison of LoID and AutoElicit across datasets in terms of magnitude, uncertainty, and performance.
A.4.2 OOD Splits
We construct out-of-distribution (OOD) train splits to simulate a training set that does not represent the entire population. For each dataset, we consider numerical features with at least two unique values and apply five shift strategies to induce distributional gaps between training and test data:
  1. 1.
    extreme_10, training on the bottom 10%
  2. 2.
    extreme_5_95, training on the bottom 5%
  3. 3.
    moderate_20_80, training on the bottom 20%;
  4. 4.
    tail_0_50, training on the 0–25% range;
  5. 5.
    tail_50_100, training on the 50–75% range.
For each split, we designate a different feature as the shift variable (cycling through available features), compute the corresponding quantile thresholds, and construct train–test masks while enforcing a minimum of 50 samples per split. Models are trained on the OOD training set and evaluated on the entire dataset to quantify performance degradation under distribution shift. Splits with insufficient class diversity (fewer than two classes in either the training or test set) are discarded. This procedure simulates realistic deployment settings in which models must generalize to regions of the feature space not observed during training.
For Heart Disease, we use the cholesterol split (moderate_20_80), training on the 20–80% quantile range and testing on extreme values, resulting in an AUC drop from 0.933 to 0.870 ($\Delta=0.063$). Liver Disease uses the alkaline phosphatase split (tail_0_50), training on the bottom 50% and testing on the top 50%, with degradation from 0.768 to 0.698 ($\Delta=0.070$). Blood Donation employs the recency split (extreme_10_90), training on donors in the bottom 10% and testing on the top 10%, experiencing the largest degradation from 0.755 to 0.579 ($\Delta=0.176$). Occupancy Detection uses the light intensity split (moderate_20_80), although this split exhibits insufficient class diversity for full evaluation. Diabetes and Pima both use the pregnancies split (extreme_10_90), training on the bottom 10% and testing on the top 10%, with degradation from 0.839 to 0.813 ($\Delta=0.027$). CreditG partitions on credit amount (extreme_5_95), training on the bottom 5% and testing on the top 5%, yielding a drop from 0.786 to 0.710 ($\Delta=0.076$). Income uses the education level split (extreme_5_95), with the steepest decline from 0.908 to 0.781 ($\Delta=0.127$). Bank Marketing splits on call duration (extreme_10_90), experiencing severe degradation from 0.888 to 0.508 ($\Delta=0.380$), representing the most extreme covariate shift in our evaluation. Finally, Jungle Chess uses the black piece strength split (tail_0_50), training on the 0–50% range and testing on 50–100%, with degradation from 0.810 to 0.712 ($\Delta=0.098$). These splits simulate realistic deployment scenarios in which models trained on one subpopulation must generalize to substantially different data distributions.
[…]
Bank 0.51 0.61 0.72 0.61 0.89 +30.0
Blood 0.58 0.63 0.67 0.62 0.75 +31.3
Credit Score 0.71 0.74 0.76 0.76 0.79 +0.0
Diabetes 0.81 0.82 0.84 0.81 0.84 +104.5
Heart 0.87 0.91 0.90 0.92 0.93 -31.8
Income 0.78 0.81 0.85 0.83 0.91 +14.9
Liver 0.70 0.68 0.69 0.70 0.77 -9.9
Occupancy 0.98 0.72 1.00 0.75 1.00 +1155.7
Jungle 0.71 0.72 0.79 0.75 0.81 +40.0
Pima 0.82 0.81 0.84 0.81 0.84 +98.1
Average 0.74 0.75 0.80 0.76 0.86 +143.3
Table 11: Comparison of LoID priors and priors with 0 means and stds in the Goldilock zone on OOD datasets. $\Delta$ Gap Closed shows the percentage point difference in gap closed between LoID and Prior = $\mathcal{N}=(0,0.45)$. OOD-LR is logistic regression trained on OOD data (lower bound), Cap is logistic regression trained on full IID data (upper bound).
A.5 LoID vs Existing Baselines
A.5.1 Prior Performance
Since LoID and AutoElicit employ similar prior elicitation strategies and exhibit more comparable performance than other baselines, Table 10 focuses on analyzing where and why each method performs better. The underlying intuition behind these differences is discussed in detail in Section 5.1.
A.5.2 Computational Efficiency
We compare the computational complexity and runtime of LoID with AutoElicit and LLMProcesses. For a dataset with $d$ features, the LoID method queries the LLM using $N_{\text{sent}}$ sentence formulations per feature and extracts the probabilities of two tokens (positive and negative). Its time complexity is:
$T_{\text{logit}}=\mathcal{O}(d\times N_{\text{sent}}\times T_{\text{LLM}}),$
where $T_{\text{LLM}}$ denotes the inference time of a single LLM query. In our experiments, $N_{\text{sent}}=10$, resulting in approximately 20 queries per feature.
AutoElicit employs an interactive multi-round elicitation process with iterative refinement, leading to a higher complexity:
$T_{\text{AE}}=\mathcal{O}(d\times N_{\text{rounds}}\times N_{\text{queries}}\times T_{\text{LLM}}),$
where $N_{\text{rounds}}\in[3,5]$ and $N_{\text{queries}}\in[5,10]$. This results in more than 28 LLM queries per feature on average.
[…]
LLMProcesses follows a fundamentally different paradigm: it generates predictions for each test instance by sampling from the LLM $N_{\text{samples}}$ times using in-context examples. For a test set of size $n_{\text{test}}$ with $n_{\text{train}}$ training examples in context, its complexity is:
$T_{\text{LLMP}}=\mathcal{O}(n_{\text{test}}\times N_{\text{samples}}\times T_{\text{LLM}}(n_{\text{train}}\cdot d)),$
where $T_{\text{LLM}}(n_{\text{train}}\cdot d)$ reflects the increased inference time caused by long prompts containing $n_{\text{train}}$ in-context examples. Unlike LoID and AutoElicit, which perform prior elicitation once per dataset, LLMProcesses scales linearly with the test set size and incurs substantially longer runtimes (15–20 hours for datasets with 100–400 test instances, compared to minutes for prior elicitation methods).
A.6 Prior Impact on Model Performance
A.6.1 LoID Priors vs. $\mathcal{N}$(0, 0.45) vs. $\mathcal{N}$(0, 1)
To isolate the contribution of the prior means inferred by LoID, we evaluate a baseline using normal priors with zero mean and standard deviation $0.45$ (the Goldilock setting). This comparison quantifies how much LoID’s learned prior means contribute to downstream AUC. As shown in Table 11, incorporating LoID-derived means yields a substantial performance gain, increasing the gap closed by approximately 154% on average. This difference is most prominent in datasets where the uninformed priors catastrophically fail: Occupancy (+1156%), Diabetes (+105%), and Pima (+98%). In these cases, the $\mathcal{N}(0,0.45)$ prior provides insufficient directional guidance, causing the Bayesian posterior to converge toward coefficients near zero and perform worse than the OOD baseline. Conversely, LoID’s informative means—despite systematic positive bias—capture feature importance magnitudes that guide the model toward the correct solution space. Even on datasets where LoID underperforms (Heart: -32%, Liver: -10%), the performance gap remains small compared to the massive gains elsewhere. This demonstrates that prior mean quality matters substantially more than perfect directional correctness when paired with appropriately calibrated uncertainty ($\sigma\approx 0.4$$0.5$).
Finally, we include an additional $\mathcal{N}(0,1)$ baseline to illustrate the improvement achieved over a more weakly informative prior that is commonly used with Bayesian inference.
A.6.2 $\mathcal{N}$(0, 1) vs. $\mathcal{U}$(-1, 1)
To understand the impact of prior distribution choice, we compare two configurations: (1) uninformative Normal priors $\mathcal{N}(0,1)$, and (2) uninformative Uniform priors $\mathcal{U}(-1,1)$.
Table 12 shows the catastrophic failure of Uniform priors (average AUC 0.42, with Occupancy at 0.02) and reveals that while Uniform priors ostensibly provide “maximum uninformativeness” by assigning equal probability across a range, they fundamentally fail as regularizers for regression coefficients because:
  1. 1.
    Lack of soft regularization: Uniform distributions do not encode the reasonable prior belief that smaller coefficient magnitudes are more plausible than larger ones. This absence of gradient-based regularization toward zero can lead to poor MCMC sampling behavior and overfitting.
  2. 2.
    Incompatibility with MCMC: The flat density of Uniform priors provides no gradient information to guide sampling algorithms, resulting in inefficient exploration and convergence issues, particularly visible in the Occupancy dataset collapse.
  3. 3.
    Prior-likelihood mismatch: For logistic regression, coefficients naturally follow a scale where values near zero are common and extreme values are rare. Uniform priors violate this structure, treating $\beta=0.1$ and $\beta=4.9$ as equally plausible a prior.
Dataset $\mathcal{N}(0,1)$ $\mathcal{U}(-1,1)$
Bank 0.61 0.35
Blood 0.63 0.59
Credit Score 0.74 0.58
Diabetes 0.82 0.32
Heart 0.91 0.84
Income 0.81 0.30
Liver 0.68 0.70
Occupancy 0.72 0.02
Jungle 0.72 0.58
Pima 0.82 0.32
Average 0.75 0.46
Table 12: Comparison of LoID priors against uninformative Normal and Uniform priors on OOD datasets.
A.7 Computational Resources
  1. 1.
    Infrastructure: Experiments were run on a workstation equipped with 2 NVIDIA RTX 6000 Ada Generation GPUs (each with 50GB VRAM), a Threadripper PRO 7985WX CPU with 64 cores (128 threads) at up to 5.37 GHz, and 256 GB of RAM.
  2. 2.
    Model Inference: We used Gemma-2-27B as our main LLM. All inference was performed using both GPUs in parallel. Each feature prompt required about 50–70 seconds to complete across 10 paraphrases, leading to approximately 60 seconds * #features per dataset.
  3. 3.
    Preprocessing: Minimal preprocessing was required and was handled efficiently on a single CPU.
  4. 4.
    Additional Resources: Exploratory analyses, including feature selection, were performed using GPT-4o via OpenRouter. Each call took 2–3 seconds and used one CPU thread.

### v2

@@
-What Language Models Know But Don’t Say: Non-Generative Prior Extraction for Generalization
+What LLMs Know but Don’t Say: Non-generative Prior Extraction for Generalization
 Abstract
@@
 Large language models contain extensive knowledge from years of research across these domains.
-We propose LoID (Logit-Informed Distributions), a deterministic method for extracting informative prior distributions for Bayesian logistic regression by directly accessing their token-level predictions.
+We propose LoID (Logit-Informed Distributions), a deterministic method for extracting informative prior distributions for Bayesian logistic and linear regression by directly accessing their token-level predictions.
 Rather than relying on generated text, we probe the model’s confidence in opposing semantic directions (positive vs.
@@
 By measuring how consistently the LLM favors one direction across diverse phrasings, we extract the strength and reliability of the model’s belief about each feature’s influence.
-We evaluate LoID on ten real-world tabular datasets under synthetic out-of-distribution (OOD) settings characterized by covariate shift, where the training data represents only a subset of the population.
-We compare our approach against (1) standard uninformative priors, (2) AutoElicit, a recent method that prompts LLMs to generate expert-style priors via text completions, (3) LLMProcesses, a method that uses LLMs to generate numerical predictions through in-context learning by formatting tabular data as natural language prompts and (4) an oracle-style upper bound derived from fitting logistic regression on the full dataset.
-We assess performance using Area Under the Curve (AUC).
-Across datasets, LoID significantly improves performance over logistic regression trained on OOD data, recovering up to 59% of the performance gap relative to the oracle model.
-LoID outperforms AutoElicit and LLMProcessesc on 8 out of 10 datasets, while providing a reproducible and computationally efficient mechanism for integrating LLM knowledge into Bayesian inference.
-What Language Models Know But Don’t Say:
-Non-Generative Prior Extraction for Generalization
-Sara Rezaeimanesh Michigan State University rezaeima@msu.edu          Mohammad M. Ghassemi Michigan State Unviersity ghassem3@msu.edu
+We evaluate LoID on 15 real-world tabular datasets under synthetic out-of-distribution (OOD) settings characterized by covariate shift.
+We compare our approach against (1) standard uninformative priors, (2) AutoElicit, a method that prompts LLMs to generate expert-style priors, (3) LLMProcesses, a method that prompts LLMs to generate predictions and (4) an oracle-style upper bound derived from fitting logistic/linear regression on an in-distribution (IID) train set.
+We assess performance using Area Under the Curve (AUC) and Mean Squared Error (MSE).
+Across datasets, LoID significantly improves performance over logistic and linear regression models trained on OOD data, recovering up to 50% of the performance gap relative to the oracle model.
 1 Introduction
@@
 (2021).
-Bayesian methods offer a principled way to encode prior knowledge and mitigate such brittleness, but manually eliciting high-quality priors at scale remains a longstanding challenge O’Hagan et al.
+Bayesian methods offer a principled way to encode prior knowledge and mitigate such brittleness, but manually eliciting high-quality priors at scale remains a longstanding challenge Mikkola et al.
 (2023).
@@
 (2025) introduced a technique to query LLMs for Gaussian priors, but its reliance on generative sampling introduces run-to-run variability.
-Parallel to this, out-of-distribution (OOD) generalization has gained increasing attention, as real-world deployment often involves data that differs from the development environment.
+While LLMs offer a promising source of priors for downstream modeling, their practical benefit depends on how well these priors generalize under distribution shift.
+OOD generalization has therefore received increasing attention, as real-world deployment often involves data that differs from the development environment.
 Meta-learning approaches like TabPFN Hollmann et al.
@@
 hidden states, logits, etc.) for tabular tasks under distribution shift.
-We address this gap with LoID (Logit-Informed Distributions), a deterministic framework that elicits informative priors for Bayesian logistic regression directly from an LLM’s token-level logits.
+We address this gap with LoID (Logit-Informed Distributions), a deterministic framework that elicits informative priors for Bayesian logistic and linear regression directly from an LLM’s token-level logits.
 Instead of relying on generated completions or examples, LoID probes the LLM with semantically paired sentences highlighting positive and negative feature-target relationships.
@@
 These priors integrate into Bayesian workflows, without any fine-tuning or in-context learning.
-We evaluate LoID on ten publicly available tabular datasets that are widely used benchmarks and have appeared in numerous prior studies, ensuring comparability with existing methods.
-We compare LoID against uninformative Bayesian inference and LLMProcesses.
-To provide an upper bound, we additionally train and evaluate a logistic regression model on the entire dataset, offering insight into performance when full population data is available.
-Across extensive out-of-distribution splits, LoID outperforms existing approaches in predictive accuracy (AUC).
-By deterministically extracting domain knowledge encoded in large language models, LoID enables robust and reproducible inference.
+We evaluate LoID on 14 publicly available tabular datasets that are widely used benchmarks and have been studied in prior work, along with one non-public medical dataset to assess performance on data that does not exist in pre-training corpora.
+We compare LoID against uninformative Bayesian inference, LLMProcesses, and AutoElicit.
+To provide an upper bound, we additionally train logistic and linear regression models on IID training sets, offering insight into performance when the true data distribution is available.
+Our experiments show that LoID improves OOD generalization across 11/15 tabular datasets, recovering 50% of the lost performance on average.
 In summary, our contributions include:
   • Logit-based deterministic prior elicitation.
-LoID is the first method to extract Bayesian priors directly from the token-level logits of large language models, eliminating the variability associated with text generation and prompt sampling.
-  • Baseline comparison across ten OOD datasets.
+LoID is the first method to extract Bayesian priors directly from the token-level logits of LLMs, eliminating the variability associated with text generation and prompt sampling.
+  • Baseline comparison across 15 OOD datasets.
 We evaluate LoID on diverse OOD tabular datasets from healthcare, finance, and social science.
@@
 Capstick et al.
-(2025) proposed AutoElicit, a method that interacts with LLMs to extract priors for Gaussian-based probabilistic models; however, this approach primarily focuses on Gaussian priors and lacks flexibility in handling other priors that may be more suitable for different data characteristics.
+(2025) proposed AutoElicit, a method that interacts with LLMs to extract priors for Gaussian-based probabilistic models.
 In a different vein, Zhong et al.
@@
 These strategies demonstrate that LLMs can inject external knowledge or expand the training distribution to mitigate generalization failure.
-Our work, LoID, complements and extends this growing body of research by proposing the first deterministic logit-based method to elicit Bayesian priors from LLMs for structured tabular tasks.
-Our method consistently improves OOD generalization across ten tabular datasets, offering a lightweight and almost deterministic integration of LLM knowledge into classical statistical pipelines.
 3 Method
 We propose LoID (Logit-Informed Distribution), a deterministic framework for extracting priors for tabular dataset features from LLMs using token-level logits, rather than relying on generated samples.
-Our goal is to derive informative priors for Bayesian logistic regression that generalize robustly in low-data settings.
+Our goal is to derive informative priors for Bayesian logistic and linear regression that generalize robustly in low-data settings.
 3.1 Problem setting
 Let $\mathcal{D}_{\text{train}}=\{(x_{i},y_{i})\}_{i=1}^{n}$ denote a training dataset, where each $x_{i}\in\mathbb{R}^{d}$ is a vector of $d$ features and $y_{i}\in\{0,1,\ldots,m\}$ is the corresponding label.
-In Bayesian logistic regression, we model the likelihood as
+In Bayesian logistic regression, we model the likelihood as:
 $p(y_{i}=t\mid x_{i},\boldsymbol{\beta})=\sigma(x_{i}^{\top}\boldsymbol{\beta}),$
 where $\sigma(\cdot)$ denotes the sigmoid function and $\boldsymbol{\beta}\in\mathbb{R}^{d}$ represents the regression coefficients.
-Bayesian inference is performed by specifying a prior distribution $p(\boldsymbol{\beta})$ and computing the posterior $p(\boldsymbol{\beta}\mid\mathcal{D}_{\text{train}}).$ In data-scarce regimes, informative priors can substantially improve generalization.
-LoID constructs such priors by leveraging domain knowledge encoded in large language models, without access to the training data.
+Bayesian inference is performed by specifying a prior distribution $p(\boldsymbol{\beta})$ and computing the posterior $p(\boldsymbol{\beta}\mid\mathcal{D}_{\text{train}}).$ In data-scarce settings, informative priors can substantially improve generalization.
+LoID constructs such priors by leveraging domain knowledge encoded in LLMs, without access to the training data.
 3.2 Querying LLM beliefs via logit differences
-For each feature $x_{j}$, we construct paired prompts that express opposing semantic claims about its effect on the target variable.
-Each prompt is carefully designed such that all relevant context appears before a sentiment-bearing token (e.g., “positive” or “negative”), ensuring that the token’s probability considers the entire context for the task, including the name of the feature and the prediction task (see examples in A.2.1).
+For each feature $f_{j}$, we construct paired prompts that express opposing semantic claims about its effect on the target variable.
+Each prompt is carefully designed such that all relevant context appears before a sentiment-bearing token (e.g., “positive” or “negative”), ensuring that the token’s probability considers the entire context for the task, including the name of the feature and the prediction task, $t$ (see examples in A.2.1).
 For example, given a prompt prefix like:
 ‘‘The impact of $f_{j}$ on $t$ is ’’
-we use the probabilities from the LLM’s softmax over its vocabulary to extract the probability of the next token being “positive” or “negative”. We define:
-$\mathcal{P}_{j}^{+}$ $=P(\texttt{``positive''}\mid\text{prefix})$
-$\mathcal{P}_{j}^{-}$ $=P(\texttt{``negative''}\mid\text{prefix})$
-where $f_{j}$ is a natural language description of feature $x_{j}$ (e.g., “age of the patient”) and $t$ describes the prediction task (e.g., “disease risk”).
-We define the logit preference score for a single prompt $p_{j}$ as:
-$\text{logit}(p_{j})=\log\left(\frac{p_{j}}{1-p_{j}}\right)$
-Where
-$p_{j}=\frac{P_{j}^{+}}{P_{j}^{+}+P_{j}^{-}}$
-The logit preference score quantifies the extent to which the model prefers the token “positive” over “negative” in a given context.
+We use the probabilities from the LLM’s softmax over its vocabulary to extract the probability of the next token being positive or negative. For the $i$th prompt prefix we define:
+$\mathcal{P}_{i}^{+}=P(\texttt{<positive\_token>}\mid\text{prefix}),\quad\mathcal{P}_{i}^{-}=P(\texttt{<negative>\_token}\mid\text{prefix})$
+We then define the logit preference score as:
+$\text{logit}_{j}(p_{i})=\log\left(\frac{p_{i}}{1-p_{i}}\right),\quad p_{i}=\frac{P_{i}^{+}}{P_{i}^{+}+P_{i}^{-}}$
+The logit preference score quantifies the extent to which the model prefers the positive token over the negative token in a given context.
 While this difference may also reflect other patterns learned during training, our hypothesis is that by constructing sentences that do not grammatically or syntactically favor either token, we reduce the influence of such confounding factors.
-As a result, the observed preference is driven primarily by the semantic relationship between the feature and the target concept.
 Finally, applying the logit transformation maps probabilities from the bounded interval (0, 1) to the unbounded real line, making the resulting scores directly appliable as coefficients in logistic regression models.
-The mean prefrence score $\mu_{j}$ for feature $x_{j}$ is given by averaging this score across a set of paraphrased prompts T{T}:
-$\mu_{j}=\frac{1}{|T|}\sum_{t\in T}\text{logit}_{t}(p_{j})$
-This value reflects the LLM’s normalized preference for a positive versus negative association with the feature across diverse prompt formulations.
-Figure 1 demonstrates how LoID extracts informative priors from model logits.
 Figure 1: LoID Pipeline: To extract informative priors, LoID uses multiple sentence templates for each feature and computes the probabilities of the tokens "positive" and "negative" as the next token.
@@
 We model each coefficient $\beta_{j}$ with a Gaussian prior: $\beta_{j}\sim\mathcal{N}(\mu_{j},\sigma_{j}^{2})$.
-The mean $\mu_{j}$ is the average logit preference score across paraphrases, which reflects the direction and strength of influence the LLM assigns to feature $x_{j}$.
-To capture uncertainty in the LLM’s belief, we define the standard deviation $\sigma_{j}$ as a scaled function of variability in the individual logit preference scores:
-$\sigma_{j}^{2}=\alpha+\gamma\cdot\text{std}(logit_{1}(p_{j}),\ldots,logit_{|T|}(p_{j}))$
-Here, $logit_{t}(p_{j})$ denotes the logit preference score for prompt $t$ and feature $j$.
-$\alpha>0$ and $\gamma>0$ are fixed scaling hyperparameters that can be tuned to optimize LoID’s performance on a certain task.
-We perform a comprehensive hyperparameter tuning in A.2.4 to find the hypermarameters that achieve the highest cumulative AUC.
-This design assigns higher variance to features for which the LLM expresses inconsistent preferences across paraphrases, thereby reflecting greater uncertainty about the feature’s influence in the model.
+The mean for feature $f_{j}$ is given by averaging this score across a set of paraphrased prompts T{T}:
+$\mu_{j}=\frac{1}{|T|}\sum_{i\in T}\text{logit}_{j}(p_{i})$
+This value reflects the LLM’s normalized preference for a positive versus negative association with the feature.
+Figure 1 demonstrates how LoID extracts informative priors from model logits.
+To capture uncertainty in the LLM’s belief, we define the standard deviation $\sigma_{j}$ as:
+$\sigma_{j}^{2}=\text{std}(logit_{1}(p_{j}),\ldots,logit_{|T|}(p_{j}))$
+Here, $logit_{i}(p_{j})$ denotes the logit preference score for the $i$th prompt and feature $j$.
+This design assigns higher variance to features for which the LLM shows inconsistent preferences across paraphrases, capturing greater uncertainty about their influence.
+We evaluated three variance computation methods and concluded that taking the standard deviation of logit preference scores as described above performs best (Section A.2.3).
 3.4 Bayesian inference with LoID priors
-With the LoID priors in place, posterior inference is performed using the No-U-Turn Sampler (NUTS), a variant of Hamiltonian Monte Carlo (HMC).
-HMC augments the parameter space with auxiliary momentum variables and explores the posterior by simulating Hamiltonian dynamics, enabling efficient, gradient-informed proposals that avoid the random-walk behavior of traditional MCMC methods.
-Specifically, HMC introduces a momentum variable $\mathbf{r}$ for the regression coefficients $\boldsymbol{\beta}$ and defines the Hamiltonian
-$\mathcal{H}(\boldsymbol{\beta},\mathbf{r})=-\log p(\boldsymbol{\beta}\mid\mathcal{D}_{\text{train}})+\frac{1}{2}\mathbf{r}^{\top}\mathbf{r}.$
-Proposals are generated by numerically simulating the resulting Hamiltonian dynamics using a leapfrog integrator, which leverages gradients of the log-posterior to propose distant yet highly probable states.
-NUTS eliminates the need to manually tune the trajectory length by adaptively extending the simulation until a U-turn condition is detected, ensuring efficient exploration while maintaining detailed balance.
-This results in faster mixing and improved effective sample sizes, particularly in high-dimensional parameter spaces.
-We run multiple NUTS chains with warm-up steps for step-size and mass-matrix adaptation, and discard burn-in samples before collecting posterior draws.
-All reported results are computed using posterior expectations estimated from these samples.
-3.5 Summary
-LoID transforms the implicit domain knowledge encoded in large language models into explicit, feature-level priors suitable for probabilistic modeling.
-Unlike prior approaches that rely on generated samples, in-context learning, or fine-tuning, LoID achieves this deterministically by directly probing token-level logits.
-This design enables reproducibility and principled integration into Bayesian models.
+With LoID priors, posterior inference is performed using a Laplace approximation, which approximates the posterior with a Gaussian centered at the maximum a posteriori (MAP) estimate:
+$\hat{\beta}=\arg\max_{\beta}\log p(\beta\mid D_{\text{train}})$ (1)
+The posterior is then locally approximated via a second-order Taylor expansion of the log-posterior:
+$p(\beta\mid D_{\text{train}})\approx\mathcal{N}(\hat{\beta},\Sigma),\quad\Sigma=\left[\nabla^{2}_{\beta}\left(-\log p(\beta\mid D_{\text{train}})\right)\big|_{\beta=\hat{\beta}}\right]^{-1}$ (2)
+This provides efficient uncertainty estimates without sampling, with LoID priors affecting both the MAP location and posterior curvature. All results are computed under this Gaussian approximation.
 3.5 Evaluation
-To assess the quality of the resulting posteriors, we compute the AUC score on the entire dataset using models trained solely on the OOD training split.
-We then compare this AUC to that of a logistic regression trained on the entire dataset.
-This comparison quantifies how much of the predictive performance gap, caused by training on an OOD distribution, is recovered by our method.
+To evaluate LoID, we construct synthetic OOD splits from real-world datasets.
+We begin by creating a standard 80–20 random split, where the training portion serves as the in-distribution (IID) training set and the held-out portion serves as the IID test set.
+To simulate distribution shift, we derive an OOD training set by selecting a subset of the IID training data and systematically excluding specific categories of categorical features or ranges of numerical features.
+Section A.5 explains the synthetic distribution shift for each dataset in detail.
+We verify the presence of distribution shift by comparing the predictive performance of a model trained on the IID training set with that of a model trained on the OOD training set, both evaluated on the same IID test set.
+A substantial degradation in performance when training on the OOD split indicates that the training data no longer adequately represents the true data distribution.
+Given this setup, we then train Bayesian models with LoID priors using the OOD training set and evaluate their performance on the IID test set.
+Our primary objective is to assess the extent to which LoID priors can mitigate the performance gap introduced by distribution shift.
 4 Experimental setup
 4.1 Datasets
-We evaluate and compare LoID’s performance on ten publicly available tabular datasets for binary and multi-class classification: The Heart Disease (Cleveland) dataset Janosi et al.
-(1989) predicts the presence of heart disease based on clinical and demographic attributes such as cholesterol level, and blood pressure.
-The Indian Liver Patient Dataset (ILPD) Ramana and Venkateswarlu (2022) classifies whether a patient has liver disease using demographic information and blood test measurements.
-The Blood Transfusion Service Center dataset Yeh (2008a) predicts whether a donor will donate blood again based on historical donation behavior.
-The Room Occupancy Detection dataset Candanedo (2016b) determines whether a room is occupied using time-series environmental sensor readings.
-The Diabetes dataset Islam et al.
-(2020) predicts the likelihood of diabetes onset based on self-reported symptoms.
-The Pima Indians Diabetes dataset Smith et al.
-(1988) predicts the presence of diabetes using diagnostic measurements such as glucose concentration and BMI.
-The Give Me Some Credit (Good Credit Score) dataset Fusion and Cukierski (2011) assesses an individual’s creditworthiness based on income, debt ratio, and delinquency history.
-The Bank Marketing dataset Moro et al.
-(2014) predicts whether a client will subscribe to a term deposit based on demographic attributes and marketing interactions.
-The Adult Income (Census Income) dataset Becker and Kohavi (1996b) predicts whether an individual’s annual income exceeds \$50K using demographic and educational features.
-Finally, the Jungle Chess (Jungle) dataset Shapiro (1983) predicts game outcomes based on structured representations of board states, providing a symbolic decision-making task.
-For each dataset, we construct synthetic OOD splits as described in Section A.4.2.
-To verify that the resulting training sets do not reflect the true data distribution, a logitic regression model is trained separately on the out-of-distribution and in-distribution training splits and evaluated on the entire dataset.
-As shown in Table 1, the model trained on the OOD train set consistently underperforms compared to the model trained on the full dataset.
-This confirms that generalization is impaired and the OOD train set does not present the full distribution of the dataset as well as the in distribution train set.
-Across all datasets, we apply the same preprocessing pipeline: categorical features are one-hot encoded, and missing values are imputed with zeros.
-For datasets with an excessive number of features—namely Give Me Some Credit, and Adult Income—we perform feature selection using GPT-4o by providing feature descriptions and prompting it to return a list of the 15 most relevant features for the prediction task.
+Head and Neck Cancer Recurrence dataset
+We compiled a binary classification dataset for Head and Neck Cancer Recurrence.
+This dataset includes comprehensive features about patient demographics, tumor characteristics, treatment regimens, and clinical staging, and predicts the likelihood of cancer recurrence following radiotherapy.
+The dataset encompasses 508 patients with head and neck cancers, capturing clinical variables including comorbidity indices (Charlson score), smoking history (pack-years), HPV status, primary tumor site localization, and detailed radiation therapy parameters such as treatment duration, prescription dose, and the critical time interval between biopsy and treatment initiation.
+This dataset presents a significant class imbalance (390 non-recurrence vs.
+118 recurrence cases), reflecting real-world clinical distributions.
+The dataset is particularly valuable for evaluating model performance under distributional shift, as we created an out-of-distribution test set focusing on elderly patients ($age\geq 75$).
+This datasets represents a clinically meaningful population that is not present in any LLM pre-training corpora.
+[^1] We will release this dataset upon publication.
+Public datasets
+We evaluate LoID on 14 publicly available tabular datasets for binary classification and regression: Adult Income Becker and Kohavi (1996), Cervical Cancer Fernandes et al.
+(2017), Chronic Kidney Disease Rubini et al.
+(2015), Bank Marketing Moro et al.
+(2014) Blood Transfusion Service Center Yeh (2008), Stroke Fedesoriano (2020), Heart Failure UCI Machine Learning Repository (2020), Diabetes Prediction Mustafatz (2023), Give Me Some Credit Kaggle (2011), Airfoil Self-Noise, Bike Sharing Fanaee-T (2013), Combined Cycle Power Plant Tfekci and Kaya (2014), Concrete Compressive Strength, Yeh (1998), and Communities and Crime Redmond (2002).
+Preprocessing
+Across all datasets, we apply the same preprocessing pipeline: categorical features are one-hot encoded, missing categorical values are imputed with mode and missing numerical values are imputed with mean.
+For datasets with an excessive number of features—namely Bank Marketing, Breast Cancer Diagnostic, Cervical Cancer, Chronic Kidney Disease, Communities and Crime, Heart Failure, Haptitis, and Head and Neck Cancer Recurrence—we perform feature selection using Claude-3.5-Sonnet by providing feature descriptions and prompting it to return a list of the 15 most relevant features for the prediction task.
 The same reduced feature set is used consistently across all experimental conditions to ensure comparability.
-Dataset AUC (IID) AUC (OOD)
-Heart Disease 0.93 0.86
-Liver Disease 0.77 0.70
-Blood Donation 0.75 0.58
-Room Occupancy 1.00 0.98
-Diabetes 0.84 0.81
-Income 0.91 0.78
-Credit 0.79 0.71
-Jungle 0.81 0.71
-Pima Diabetes 0.84 0.81
-Bank Marketing 0.89 0.51
-Table 1: AUC comparison of logistic regression on IID vs. OOD test splits. Lower performance on OOD indicates distribution shift.
-We designate the Heart Disease, Blood Donation, and Diabetes datasets as development datasets for evaluating prompt structures and hyperparameter tuning.
-The remaining five datasets are reserved for testing the generalization performance of the final LoID method.
-See (A.2.4) for details on the hyperparameter search process and value ranges explored.
-Dataset OOD-LR AutoElicit (baseline) LLMProc (baseline) LoID (our method) Cap (upper bound) Gap Closed%
-Heart Disease 0.87 0.92 0.86 0.90 0.93 +50
-Liver Disease 0.70 0.36 0.60 0.69 0.77 -19
-Blood Donation 0.58 0.71 0.63 0.67 0.75 +52
-Room Occupancy 0.98 0.97 0.80 1.00 1.00 +78
-Diabetes 0.81 0.83 0.70 0.84 0.84 +99
-Good Credit Score 0.71 0.71 0.30 0.76 0.79 +60
-Income 0.78 0.85 0.51 0.87 0.91 +57
-Pima Diabetes 0.82 0.83 0.74 0.84 0.84 +88
-Bank Marketing 0.51 0.51 0.51 0.72 0.89 +56
-Jungle Chess 0.71 0.78 0.62 0.79 0.81 +70
-Table 2: Comparison of AUC scores across datasets.
-“OOD-LR” reports the performance of logistic regression trained on an OOD training set.
-LoID results are obtained using priors from Gemma-2-27B, with $\alpha=0.2$, $\gamma=2.0$ and $N_{sent}=10$.
-Among Uniformative, AutoElicit, LLMProcesses, and LoID, the highest AUC for each dataset is shown in bold.
-The Cap condition serves as an oracle-style upper bound.
-All values are rounded to two decimal places.
-In some cases, LoID and Cap appear to have similar performance; however, the gap is not fully closed.
-This is due to rounding.
-For example, on Diabetes, LoID achieves 0.839 while Cap achieves 0.842.
+We apply z-score normalization to the features for AutoElicit and LoID, but not for LLMProcesses, following the authors’ recommendations.
+Development datasets
+We designate the Stroke, Blood Donation, Heart Failure, and Adult Income datasets as development datasets for choosing an LLM and hyperparameter tuning A.2.
+The remaining 11 datasets are reserved for testing the generalization performance of the final LoID method.
+4.2 Creating distribution shifts
+For each dataset, we construct synthetic OOD splits as described in Section A.5.
+As shown in Table 1 in the P_IID (in-distribution train set performance), and P_OOD (out-of-distribution train set performance) columns, the model trained on the OOD train set noticeably underperforms compared to the model trained on the IID train set.
+This confirms that generalization is impaired and the OOD train set does not represent the full distribution of the dataset as well as the IID train set.
+Dataset P_OOD P_IID N(0,1) LoID LLMProc AutoEl Gap (%)
+Cap our method Requeima et al. (2024) Capstick et al. (2025)
+Classification (AUC)
+Head & Neck Cancer 0.542 0.998 0.567 0.696 0.540 0.609 33.8
+Adult Income 0.414 0.809 0.51 0.745 0.603 0.510 83.8
+Cervical Cancer 0.440 0.870 0.448 0.681 0.658 0.570 56.0
+Kidney Disease 0.822 0.993 0.715 0.901 0.590 0.320 46.2
+Diabetes 0.883 0.963 0.853 0.906 0.502 0.864 28.8
+Stroke 0.608 0.842 0.618 0.694 0.500 0.670 36.5
+Heart Failure 0.746 0.849 0.742 0.753 0.770 0.748 6.8
+Bank Marketing 0.780 0.878 0.780 0.652 0.498 0.692 -130.6
+Blood Donation 0.572 0.748 0.572 0.663 0.483 0.556 51.7
+Good Credit Score 0.710 0.790 0.710 0.760 0.700 0.710 62.5
+Regression (MSE)
+Concrete Strength 469.20 98.14 165.72 120.59 136.57 1974.11 93.9
+Communities & Crime 0.17 0.01 0.073 0.07 0.15 0.30 62.5
+Airfoil Noise 77.16 23.69 51.11 60.51 47.80 62.21 68.9
+Cycle Power 962.94 21.24 23.58 509.92 462.89 47.14 0.1
+Bike Sharing 77858 19030 71890 78351 75650 71405 -1.8
+Table 1: Performance comparison across 15 datasets.
+Gap Closed represents the percentage of the OOD performance gap recovered by our method, LoID.
+The percentage closed in calculated using: $\frac{|LoID-P\_OOD|}{|P\_IID-P\_OOD|}\times 100$
 4.3 LLM-based prior extraction
-We use the Gemma-2-27B, Gemma-3-27B, llama-3-8B, and GPT-2 large language models to extract logit-based priors as described in Section 3.2.
+We use Gemma-2-27B, Gemma-3-27B, llama-3-8B, and GPT-2 to extract logit-based priors as described in Section 3.2.
 We use these models in zero-shot mode without any fine-tuning or in-context examples.
@@
 For each feature, we issue 10 prompts to extract token-level log-probabilities for the positive and negative completions.
-The number of prompts was selected via hyperparameter tuning (Section A.2.4).
-The resulting feature-wise prior means and variances are then used to define informative Gaussian priors in a Bayesian logistic regression model.
+The number of prompts is selected via hyperparameter tuning (Section A.2.1).
+The resulting feature-wise prior means and variances are then used to define informative Gaussian priors in Bayesian logistic and linear regression models.
 Posterior inference is performed using PyMC with a Laplace approximation.
-We compare LoID to the following baselines:
-  • OOD-LR: Logistic regression model trained on OOD train set.
-  • AutoElicit: LLM-generated expert-style priors via sampled completions.
+We compare LoID to the following baselines.
+  • P_OOD: Logistic/Linear regression model trained on OOD train set.
+  • AutoEl (AutoElicit): LLM-generated expert-style priors via sampled completions.
   • LLMProc (LLMProcesses): In-context learning approach that generates numerical predictions by formatting tabular data as natural language prompts.
 For fair comparison, we use the OOD training examples as few shot examples for the LLM.
-  • Cap: An oracle-style upper bound achieved by training a logistic regression model on the entire dataset, simulating access to the full population.
-We evaluate each method using AUC score obtained from training on the ood train set and testing on the entire dataset.
-5 Results
-Table 2 summarizes the performance of each method across ten tabular datasets.
-Based on our experiments in Section A.2.2, Gemma-2-27B achieves the highest cumulative AUC among the evaluated LLMs.
-Therefore, Table 2 reports results obtained using this model.
-5.1 AUC Comparison
-LoID outperforms logistic regression trained on the OOD training set for all datasets except Liver Disease.
-Notably, on the Diabetes and Pima Diabetes datasets, LoID almost completely closes the performance gap induced by OOD training, indicating that it can successfully infer the underlying data distribution.
-On average, LoID closes 59% of the performance gap introduced by the OOD training setting.
-LLMProcesses often underperforms even compared to uninformative priors.
-Since this approach does not explicitly produce priors, it is difficult to determine the exact cause of its poor performance.
-One possible explanation is its complete reliance on LLMs without leveraging traditional machine learning models, which are often more robust at pattern extraction in structured data.
-LoID also achieves superior performance on 8 out of 10 datasets, achieving an average improvement of 99.7% in gap closed when it outperforms AutoElicit.
-To further understand these performance differences, we analyze the elicited prior distributions produced by both methods across selected winning and losing datasets.
-Section A.5.1 presents this comparison.
-Finding 1: Non-Monotonic Relationship Between Prior Uncertainty and Performance
-  • Optimal range ($\sigma\approx 0.4$$0.5$): Achieves best performance on 7 datasets: Liver, Income, Bank, Creditg, Diabetes, Pima, Occupancy
-  • Too weak ($\sigma\gtrsim 0.8$): Underperforms AutoElicit: Blood, Heart
-This shows that moderately weak priors ($\sigma\approx 0.4$$0.5$) provide sufficient regularization while allowing the Bayesian posterior to correct systematic biases in the elicited means through the data likelihood.
-Excessively weak priors ($\sigma\gtrsim 0.8$) fail to leverage LLM knowledge effectively, approaching uninformed priors.
-LoID excels when Prior uncertainty is moderate, datasets have sufficient samples to correct sign biases, and AutoElicit’s tight priors impose incorrect strong beliefs.
-Finding 3: Superior Mean Estimation Despite Directional Bias
-While LoID exhibits systematic positive bias in sign coverage (0–6% negative priors vs.
-AutoElicit’s 20–50%), the absolute quality of the mean estimates appears superior.
-Evidence from performance gaps suggests:
+  • P_IID: Performance obtained by training a logistic/linear regression model on the in-distribution (IID) data.
+This represents the oracle benchmark where the model has full access to the training distribution, serving as a reference point for evaluating OOD generalization.
+  • N(0,1): An uninformed baseline using standard normal $\mathcal{N}(0,1)$ priors for all feature coefficients.
+This represents a weakly informative Bayesian approach without incorporating domain knowledge, serving as a reference point to quantify the information gain provided by LLM-derived priors.
+While a uniform prior would be maximally uninformative, our experiments in Section A.2.4 show that normal priors are more compatible with the MCMC pipeline and provide a stricter baseline.
+For AutoElicit and LLMProcesses, we adopt the hyperparameters recommended by the original authors to the extent possible in order to ensure a fair comparison.
+A comprehensive list of the hyperparameters used to reproduce these baselines is provided in Section A.1.
+5 Results and discussion
+We evaluate LoID across 15 datasets spanning diverse domains to understand when LoID is able to close the performance gap caused by covariate shift.
+Our results in Table 1 reveal a systematic pattern: LoID excels when feature-target relationships are semantically meaningful and generalizable, but struggles when relationships are contextual, behavioral, or dataset-specific.
+5.1 When does LoID help?
+Medical and demographic tasks with stable relationships (83.8% gap closed on average).
+LoID achieves the strongest gains on datasets where feature-target relationships reflect generalizable biomedical or socioeconomic knowledge:
+  • Adult Income (83.8% gap): Age, education, and occupation have well-known effects on income that generalize across demographic shifts
+  • Concrete Strength (93.9% gap): Physical properties (cement, age, water) have deterministic relationships with compressive strength
+  • Chronic Kidney Disease (46.2% gap): Clinical markers (hemoglobin, albumin) have established diagnostic value
+In these domains, the LLM’s world knowledge captures stable causal mechanisms that remain valid under distribution shift.
+Physical systems with known causal structure.
+For regression tasks where features have clear physical interpretations (Concrete Strength: 93.9% gap, Airfoil Noise: 68.9% gap, Communities and Crime: 62.5% gap), LoID leverages the LLM’s understanding of domain physics and social science to regularize predictions under shift.
+Figure 2: LLM prior alignment across success and failure modes.
+Top: Well-aligned priors for Blood Donation and Chronic Kidney Disease show strong agreement between LLM beliefs and actual correlations.
+Bottom: Combined Cycle Power exhibits directionally correct but magnitude-misspecified priors, while Bank Marketing shows absent priors (near-zero) for context-dependent features.
+5.2 When does LoID hurt?
+Behavioral and contextual tasks with dataset-specific patterns. LoID fails on three datasets where relationships are inherently local or behavioral:
+  • Bank Marketing ($130.6\%$ gap): Customer banking behavior (call duration, previous contacts) varies wildly across campaigns.
+  • Bike Sharing ($1.8\%$ gap): Demand is driven by local weather, events, and infrastructure, relationships that are city-specific and temporal
+  • Combined Cycle Power ($0.1\%$ gap): Plant efficiency depends on equipment-specific calibrations and operational conditions that LLMs cannot infer from generic turbine descriptions
+In these cases, the LLM’s priors encode generic relationships (e.g., “higher temperature decreases power output”) that may be directionally correct but quantitatively misspecified for the specific plant, city, or campaign.
+This introduces harmful bias that degrades performance.
+Understanding failure modes through prior analysis. Examining the extracted priors reveals two distinct failure modes:
   1. 1.
-    Magnitude accuracy over sign accuracy: On datasets where LoID wins, the improvement over AutoElicit indicates that the LLM-elicited means—despite sign bias—capture feature importance better than AutoElicit’s conversational approach.
+    Overconfident but wrong priors destroy beneficial regularization.
+For Combined Cycle Power, the uninformed $\mathcal{N}(0,1)$ baseline closes 99.8% of the OOD gap (MSE: 23.58 vs.
+P_OOD: 962.94), demonstrating that simple Bayesian regularization is highly effective.
+However, LoID closes only 0.2% (MSE: 961.04), because the LLM’s equipment-agnostic priors (e.g., “higher ambient temperature decreases turbine efficiency”) override the beneficial regularization with domain knowledge that’s directionally plausible but quantitatively misspecified for this specific power plant and severe temperature shift (train: 1.8–13.5°C, test: 13.5–37.1°C).
+The LLM knows generic thermodynamics but cannot account for plant-specific calibrations or extreme extrapolation regimes.
   2. 2.
-    Comparison to uninformed priors: LoID achieves  145% better gap closed vs.
-uninformed priors (see full comparison in A.6.1), demonstrating that the mean estimates contain genuine signal beyond just the regularization effect of uncertainty.
-  3. 3.
-    Interaction with uncertainty: The combination of (a) approximately correct magnitudes in means and (b) moderate uncertainty allows the Bayesian posterior to correct sign errors while preserving magnitude information.
-In contrast, AutoElicit’s tight uncertainties amplify errors in both magnitude and direction.
-5.2 Computational Efficiency.
-[…]
-6 Discussion
-The success of our LoID method challenges the conventional assumption that more accurate prior means necessarily lead to better performance.
-AutoElicit provides strong, confident priors ($\sigma\approx 0.05$$0.12$) with $\sim$40–50% correct negative signs.
-However, the incorrect confident priors harm inference by overpowering the data likelihood.
-On the other hand, LoID provides weaker and uncertain priors ($\sigma\approx 0.4$$0.5$).
-This allows the weaker means to be overridden by data, providing soft regularization: the priors gently guide inference toward reasonable parameter spaces without rigidly enforcing incorrect beliefs.
-The result is that even when LoID’s directional beliefs are wrong, the posterior can recover because the prior is not insisting on its incorrectness.
-However, the benefits of this approach diminish when priors become too weak ($\sigma>0.8$).
+    Context-dependent features produce uninformative priors.
+For Bank Marketing and Bike Sharing, LLMs have no basis for judging how campaign-specific variables (e.g., “days since last contact”) or city-specific infrastructure affects outcomes.
+The extracted priors are either near-zero (indicating maximal uncertainty) or directionally inconsistent, providing no signal over uninformed regularization.
+Unlike Combined Cycle Power, where the problem is wrong priors, here the problem is absent priors—the LLM simply lacks relevant knowledge about these context-dependent domains.
+Figure 2 illustrates four representative cases that reveal when and why LLM priors succeed or fail.
+5.3 Comparison to baselines
+LoID vs.
+LLMProcesses.
+LLMProcesses attempts to directly predict outcomes by prompting the LLM with feature values.
+While this works for more common tasks, it fails on real-world data: LLMProcesses achieves near-random performance (0.500 AUC) on 4/9 classification tasks (Stroke, Diabetes, Heart Failure, HNC).
+LoID outperforms LLMProcesses on 12/14 datasets (excluding Bank Marketing where both fail).
+This demonstrates that extracting priors is more effective than direct prediction—the LLM knows the relationships but cannot directly predict outcomes from sparse feature descriptions.
+LoID vs.
+AutoElicit.
+AutoElicit elicits structured expert priors by prompting the LLM multiple times and aggregating responses to reduce variability.
+It performs well on tasks involving physical systems (Bike Sharing: 71405 MSE; Combined Cycle Power: 47.14 MSE), where aggregating numeric estimates helps calibrate prior ranges.
+However, for tasks requiring semantic reasoning about feature importance (Adult Income: 0.510 AUC; Concrete Strength: 1974.11 MSE), free-form elicitation may struggle to capture directional relationships as effectively as token-level probability extraction.
+Moreover, LLMs tend to produce overconfident priors when explicitly prompted, which can limit the model’s ability to adapt to data.
+In contrast, LoID derives priors directly from token probabilities, leading to more calibrated uncertainty.
+As a result, LoID outperforms AutoElicit on 12/14 datasets.
+The N(0,1) baseline reveals when priors help.
+Comparing LoID to the uninformed $\mathcal{N}(0,1)$ baseline isolates the value of LLM knowledge.
+On 11/15 tasks, LoID substantially outperforms $\mathcal{N}(0,1)$, confirming that LLM priors are informative.
 6 Limitations and future work
 While our results demonstrate the effectiveness of LoID, several limitations should be noted, which also point toward directions for future work.
-Semantic granularity.
-Our current approach probes binary directional relationships (positive vs.
-negative).
-Extending this framework to multi-token or scalar outputs could capture more nuanced associations, such as non-monotonic effects.
-Hyperparameter selection:
-Although we use fixed hyperparameters ($\alpha=0.2$, $\gamma=2.0$) across all test datasets, optimal values vary by task.
-Developing automated calibration methods or meta-learning approaches for selecting these parameters would improve out-of-the-box applicability.
-LLM access requirements:
+  • LoID is most useful for dataset with meaningful semantic features, limiting its usage for de-identified datasets.
   • LoID requires access to token-level logits, limiting compatibility with closed-source APIs.
 7 Conclusion
-In this work, we introduce LoID, a deterministic method for extracting priors from LLMs.
-We show that priors elicited with LoID substantially mitigate the performance degradation caused by training on small, out-of-distribution datasets, recovering over 59% of the lost performance.
-Our comparative analysis with AutoElicit demonstrates that the effectiveness of the LoID approach stems from honest uncertainty quantification, and correct mean magnitude estimation for logistic regression coefficients.
-By maintaining moderate prior uncertainty ($\sigma\approx 0.4$$0.5$), LoID provides soft regularization while allowing the Bayesian posterior to correct systematic biases through the data likelihood.
-As a result, LoID outperforms AutoElicit on 80% of benchmark datasets, achieving an average improvement of 99.7% in gap closed when it wins.
-Against LLMProcesses, LoID achieves superior performance on all 10 datasets, with particularly dramatic gains on datasets where LLMProcesses’s end-to-end LLM approach fails to leverage structured data patterns.
+We introduced LoID, a method for extracting Bayesian priors from large language models through token-level probability distributions.
+By probing LLMs for beliefs about feature–target relationships and converting these into structured priors, LoID leverages embedded world knowledge to improve predictive performance under distribution shift.
+Across 15 datasets, we observe a consistent pattern: LoID performs strongly when feature–target relationships are semantically stable and generalizable, closing 83.8% of the performance gap induced by out-of-distribution training in medical and demographic settings, but degrades when these relationships are context-dependent or dataset-specific, as in behavioral tasks such as Bank Marketing.
+Overall, LoID demonstrates that extracting distributional priors is more robust than direct prediction (compared to LLMProcesses) and more reliable than free-form elicitation (compared to AutoElicit), outperforming both baselines on 12 out of 15 datasets.
 Appendix A Appendix
-A.1 Reproducibility
-Code Availability
-To ensure reproducibility, we will release our full codebase, including detailed instructions for running the experiments and reproducing all reported results.
-We also provide the the posterior files generated by our experiments, as well as the scripts used to generate all figures and tables in the paper.
-Data Availability
-All datasets used in this study are publicly available through open-access repositories.
-Further details and references are provided in Section A.4.
-All preprocessing steps required to obtain the final train/test splits used in our experiments, along with the preprocessed data, are available in our code release.
-Results Variability
-To assess the stability and determinism of our approach, all experiments were repeated five times.
-The resulting 95% confidence intervals for AUC were effectively zero (0.000), indicating that our method yields highly consistent results across runs.
+A.1 Baseline hyperparameters
+This section details the hyperparameters used for the baseline methods compared in our experiments. Table 2 provides a comprehensive overview of all configuration settings.
+Category LLMProcesses AutoElicit LoID
+Model Configuration
+Model Gemma-2-27B-it Gemma-2-27B-it Gemma-2-27B-it
+Temperature 1.0 1.0 N/A (no sampling)
+Top-p 1.0 0.9 N/A
+Max length 20 tokens 8192 tokens N/A
+Prior/Prediction Extraction
+Method Monte Carlo sampling Multi-expert prompting Token probability
+Samples per point 50 500 (prior predictive) 10 templates
+Aggregation Median MCMC mixture Mean and Variance of log-probs
+Templates/Roles N/A 100 (10×10 role pairs) 10 sentence templates
+Data Configuration
+Training samples 50 (30 for HNC) 25 N/A (no in-context)
+Prompt ordering Distance-based N/A N/A
+Reproducibility
+Random seed N/A 42 42
+Repetitions 5 5 5
+Table 2: Hyperparameter configurations for AutoElicit, LLMProcesses, and LoID.
 A.2 Hyperparameter optimization
-We used three datasets—Blood Donation, Heart Disease, and Diabetes—as development datasets to tune hyperparameters based on AUC score.
+We used four datasets—Stroke, Blood Donation, Hear Failure, and Adult Income—as development datasets to tune hyperparameters based on AUC and MSE.
 A.2.1 Number of prompt templates
-We tested our method with 1–20 prompt templates analyzed how the number of templates affects LoID’s performance. The following are some examples of our prompt templates:
+We tested our method with 2–20 prompt templates and analyzed how the number of templates affects LoID’s performance. The following are some examples of our prompt templates:
   1. 1.
@@
     The correlation between and is …
-As shown in Figure 2, performance gains plateau after approximately 10 sentences.
-Around this point, performance also becomes more stable, while incorporating additional sentences incurs increased computational cost.
-Accordingly, we use 10 sentences in all experiments.
-Figure 2: Hyperparameter ablation results demonstrating that prior uncertainty (beta) is the critical factor for performance, with F-statistics of 346.7 (blood), 47.3 (heart), and 583.2 (diabetes), while variance sensitivity (alpha) and number of sentence formulations (num_sentences) show no significant effect.
+Figure 3 reveals the optimal template count for prior extraction.
+Performance peaks at 10 templates across most datasets, balancing signal aggregation against overfitting.
+Beyond 10 templates, performance degrades sharply (e.g., Stroke drops from 0.69 to 0.39 at 15 templates), suggesting that excessive prompting amplifies misspecified priors rather than refining them.
+While coefficient of variation decreases with more templates (indicating more stable uncertainty estimates), this stability comes at the cost of performance when templates exceed 10.
+Notably, heterogeneity reduction plateaus after 10 templates, yet performance continues to degrade, demonstrating that template averaging helps only when the underlying priors are well-specified.
+For datasets where LLM knowledge is weak (e.g., Blood), even 2 templates outperform 15+, as additional prompting merely reinforces uninformative/incorrect beliefs.
+This pattern supports our recommendation of 10 templates as the optimal trade-off between variance reduction and avoiding overconfident misspecification.
+Figure 3: Template count sensitivity analysis across four datasets.
+Top two rows: Performance (Test AUC) versus number of sentence templates, comparing LoID to out-of-distribution baseline (P_OOD).
+Performance peaks at 10 templates for most datasets, with degradation at 15+ templates suggesting overfitting to misspecified priors.
+Bottom two rows: Prior heterogeneity measured by coefficient of variation (CV) of extracted prior standard deviations.
+Higher CV indicates greater variance in uncertainty across features.
+Heterogeneity decreases with more templates as averaging stabilizes estimates, but excessive templates (15+) harm performance without further reducing uncertainty.
 A.2.2 Model selection
-We compared the performances of LLMs–GPT-2, LLaMA 3.1 8B, Gemma-3-27B, and Gemma-2-27B in terms of AUC. The best averge performance was achieved with Gemma-2-27B, as shown in Table 3.
-dataset blood diabetes heart average
-model
-gemma-3 0.65 0.84 0.89 0.79
-gemma-2 0.67 0.84 0.90 0.80
-llama-3-8b 0.56 0.80 0.85 0.73
-gpt-2 0.53 0.82 0.84 0.71
-Table 3: Model comparison across datasets using $\alpha=0.2$, $\gamma=2.0$, and 10 sentence formulations.
+We compared the performances of LLMs–GPT-2 (GPT), LLaMA-3.1-8B (L-3), Gemma-3-27B-Instruct (G-3), and Gemma-2-27B-Instruct (G-2) in terms of AUC.
+The best average performance was achieved with Gemma-2-27B, as shown in Table 3.
+Dataset G-2 G-3 L-3 GPT-2
+Stroke 0.694 0.690 0.624 0.712
+Blood 0.663 0.522 0.609 0.579
+Heart Failure 0.753 0.767 0.737 0.585
+Adult 0.745 0.562 0.550 0.370
+Average 0.714 0.635 0.630 0.562
+(a) Model ablation.
+Dataset M-1 M-2 M-3
+Stroke 0.694 0.686 0.690
+Blood 0.663 0.576 0.575
+Heart Failure 0.753 0.708 0.743
+Adult 0.745 0.674 0.659
+Average 0.714 0.661 0.667
+(b) Variance method ablation using Gemma-2-27B.
+Table 3: Choosing the best model and variance estimate using four development datasets (Stroke, Blood Donation, Heart Failure, Adult Income).
 A.2.3 Variance estimation formula
@@
 We explored two methods for computing $\sigma_{j}$ from LLM token probabilities, each capturing different aspects of elicitation uncertainty.
-Method 1: Variance of Logit Differences
-This method uses the variance across multiple sentence formulations as a measure of prior uncertainty:
-$\sigma_{j}^{2}=\alpha+\gamma\cdot\text{std}(logit_{1}(p_{j}),\ldots,logit_{|T|}(p_{j}))$
+The best average performance was achieved with Method 1, as shown in Table 3.
+Method 1: variance of normalized logit probabilities
+In this method, $p_{j}$ reflects the model’s preference between logits.
+Applying a log transform maps probabilities from $[0,1]$ to the real line, aligning with the unbounded nature of regression coefficients.
+Variance across paraphrased inputs captures inconsistency in LLM responses, indicating greater epistemic uncertainty.
+$\sigma_{j}^{2}=\text{std}(logit_{1}(p_{j}),\ldots,logit_{|T|}(p_{j}))$
 where
@@
 $p_{j}=\frac{P_{j}^{+}}{P_{j}^{+}+P_{j}^{-}}$
-Rationale: Higher variance across sentence formulations indicates inconsistent LLM responses, suggesting greater epistemic uncertainty.
-The hyperparameters $\alpha$ and $\gamma$ control the base uncertainty level and sensitivity to variance, respectively.
 Method 2: entropy-based uncertainty
-This method uses Shannon entropy to measure the uncertainty in the LLM’s token probability distribution:
-$\sigma_{j}=\beta\cdot\frac{1}{N_{\text{sent}}}\sum_{i=1}^{N_{\text{sent}}}H(p_{\text{pos}}^{(i)},p_{\text{neg}}^{(i)})$
-Rationale: Higher entropy (closer to $\log 2\approx 0.693$) indicates the LLM is uncertain between “positive” and “negative” tokens, reflecting ambiguity about the feature’s effect direction.
+This method uses Shannon entropy to measure the uncertainty in the LLM’s token probability distribution.
+Higher entropy (closer to $\log 2\approx 0.693$) indicates the LLM is uncertain between the positive and negative tokens, reflecting ambiguity about the feature’s effect direction.
 Lower entropy indicates confident predictions.
-The scaling factor $\beta$ maps entropy to prior standard deviation.
-Results
-Table 4 compares the variance estimation methods described above.
-While the overall differences in performance are small, the variance of logit differences consistently performs slightly better across all three datasets.
-Accordingly, we adopt this method for LoID.
-Dataset variance Entropy $\Delta$
-Blood 0.67 0.66 +0.01
-Heart 0.90 0.89 +0.01
-Diabetes 0.84 0.83 +0.01
-Average 0.80 0.79 +0.01
-Table 4: Comparison of variance of logit differences (used by LoID) vs. entropy-based uncertainty for uncertainty estimation.
-A.2.4 Tuning $\alpha$, $\gamma$, and $N_{\text{sent}}$
-As detailed in Section 3.3, we use the following formula to calculate standard deviation:
-$\sigma_{j}^{2}=\alpha+\gamma\cdot\text{std}(logit_{1}(p_{j}),\ldots,logit_{|T|}(p_{j}))$
-We searched over $\alpha\in\{0.1,0.2,0.3,0.5\}$, $\gamma\in\{1,2,3,4\}$, and $N_{\text{sent}}\in\{5,10,15,20\}$.
-Table 5 reports the hyperparameter settings that achieve the best performance for each evaluation dataset.
-As shown in the table, no single hyperparameter configuration generalizes across all datasets; instead, each dataset requires task-specific tuning to achieve optimal performance.
-Among the evaluated $\alpha$$\gamma$ combinations, $\alpha=0.2$, $\gamma=2.0$, and $N_{\text{sent}}=10$ yield the best average performance.
-For simplicity, we adopt these values for all remaining experiments.
-Dataset Best AUC $\alpha$ $\beta$ $N_{\text{sent}}$
-Blood 0.6704 0.2 4.0 10
-Heart 0.8987 0.1 2.0 20
-Diabetes 0.8392 0.5 1.0 5
-Overall (avg) 0.8027 0.2 2.0 10
-Table 5: Best hyperparameter configurations identified for each dataset in the ablation study.
-Figure 2 illustrates the relationship between AUC and the hyperparameters $\alpha$, $\gamma$, and the number of sentences.
-Among these, $\gamma$ has the most pronounced effect on performance, with its influence on AUC appearing approximately linear.
-For the Blood Donation dataset, AUC increases as $\gamma$ increases, whereas for the Heart Disease and Diabetes datasets, AUC decreases with larger values of $\gamma$.
-These trends indicate that prior certainty plays a pivotal role in model performance.
-In contrast, we observe no consistent relationship between $\alpha$, the number of sentences, and AUC across datasets, suggesting that these factors have a limited impact on performance.
-Consistent with the previous findings, analysis of variance (ANOVA) reveals that $\gamma$ is the only statistically significant predictor of out-of-distribution performance across all datasets (Table 6).
-The baseline uncertainty parameter $\gamma$ exhibits strong effects ($p<0.0001$).
-In contrast, both $\alpha$ and $N_{\text{sent}}$ show no significant impact ($p>0.05$) on any dataset.
-Dataset Parameter F-statistic p-value
-Blood $\alpha$ 0.231 0.8743
-$\gamma$ 346.743 <0.0001<0.0001
-$N_{\text{sent}}$ 0.436 0.7280
-Heart $\alpha$ 0.360 0.7820
-$\gamma$ 47.269 <0.0001<0.0001
-$N_{\text{sent}}$ 1.174 0.3285
-Diabetes $\alpha$ 1.162 0.3331
-$\gamma$ 583.151 <0.0001<0.0001
-$N_{\text{sent}}$ 0.083 0.9691
-Table 6: ANOVA results testing hyperparameter effects on out-of-distribution AUC. Only $\gamma$ (baseline prior uncertainty) shows significant effects across all datasets.
+$\sigma_{j}=\frac{1}{N_{\text{sent}}}\sum_{i=1}^{N_{\text{sent}}}H({P_{j}^{-}},P_{j}^{+})$
+Method 3: standard deviation of log-probability differences
+This method uses the difference of log-probabilities (equivalent to the log-ratio $\log(\frac{p_{\text{pos}}}{p_{\text{neg}}})$) without normalization, preserving information about the absolute probability mass assigned to both tokens.
+When templates produce inconsistent log-probability differences, this indicates epistemic uncertainty about both the direction and confidence of the feature’s effect.
+Unlike the normalized logit approach, this method is sensitive to cases where the LLM assigns low absolute probability to both tokens (indicating high uncertainty) versus high probability to one token (indicating confidence).
+$\sigma_{j}^{2}=\text{std}(diff_{1}(p_{j}),\ldots,diff_{|T|}(p_{j}))$
+Where
+$diff(p_{j})=[\log(P_{j}^{+})-\log(P_{j}^{-})]$
+A.2.4 Uninformative baseline: $\mathcal{N}$(0, 1) vs. $\mathcal{U}$(-1, 1)
+We include an additional $\mathcal{N}(0,1)$ baseline to illustrate the improvement achieved over a more weakly informative prior that is commonly used with Bayesian inference.
+To understand the impact of prior distribution choice, we compare two configurations: (1) uninformative Normal priors $\mathcal{N}(0,1)$, and (2) uninformative Uniform priors $\mathcal{U}(-1,1)$.
+Table 4 reveals that while Uniform priors ostensibly provide “maximum uninformativeness” by assigning equal probability across a range, they fail as regularizers for regression coefficients because:
+  1. 1.
+    Lack of soft regularization: Uniform distributions do not encode the reasonable prior belief that smaller coefficient magnitudes are more plausible than larger ones.
+This absence of gradient-based regularization toward zero can lead to poor MCMC sampling behavior and overfitting.
+  2. 2.
+    Incompatibility with MCMC: The flat density of Uniform priors provides no gradient information to guide sampling algorithms, resulting in inefficient exploration and convergence issues.
+Dataset N(0,1) U(-1,1) $\Delta$
+Stroke 0.694 0.448 -0.246
+Blood 0.583 0.575 -0.008
+Heart Failure 0.752 0.724 -0.028
+Adult 0.745 0.761 +0.016
+Mean 0.694 0.627 -0.066
+Table 4: Comparison of Normal vs. Uniform uninformed priors for Bayesian logistic/linear regression.
 A.3 Additional experiments
@@
 Table 5 presents examples of these features and their corresponding synonyms.
-Dataset Original Feature Synonym Set 1 Synonym Set 2
-Blood recency Time since last donation Days elapsed since previous donation
-frequency Number of donations made Donation count
-Heart Age Patient age Years of life
-Cholesterol Serum cholesterol level Blood cholesterol concentration
-MaxHR Peak heart rate Highest heart rate reached
-ExerciseAngina Angina during exercise Exercise-provoked chest pain
-Diabetes Glucose Blood glucose level Plasma glucose concentration
-BMI Body mass index Weight-to-height ratio
-DiabetesPedigree Family diabetes history score Genetic diabetes risk
-Table 7: Representative examples of feature name variations tested in synonym ablation study (Synonym Set 0 = original feature names).
-Results in Table 8 reveal minimal performance variation across three distinct feature naming schemes, with AUC ranges of 0.007, 0.002, and 0.001 for blood, heart, and diabetes datasets respectively.
-These negligible differences demonstrate that our elicitation method is robust to superficial linguistic variations in feature descriptions.
-This stability is significant because it indicates the LLM is capturing genuine semantic understanding of the features’ predictive relationships rather than relying on brittle pattern matching or memorized associations with specific phrasings.
-Dataset Syn0 Syn1 Syn2 Range
-Bank 0.692 0.692 0.696 0.004
-Blood 0.657 0.654 0.655 0.004
-Credit Score 0.755 0.755 0.755 0.000
-Diabetes 0.836 0.835 0.835 0.000
-Heart 0.887 0.887 0.887 0.000
-Income 0.844 0.842 0.844 0.002
-Liver 0.667 0.666 0.667 0.001
-Occupancy 0.988 0.988 0.988 0.001
-Jungle 0.79 0.778 0.778 0.012
-Pima 0.836 0.836 0.837 0.000
-Mean Range 0.0013
-Table 8: Synonym Ablation Study: Robustness to feature name variations across all datasets. Three synonym sets are tested.
-A.3.2 Feature Description
-We analyzed the impact of adding further description to the features and the target could help the LLM arrive at better estimates of the prior distributions.
-Table 9 reveals minimal and inconsistent impact from including feature names alongside descriptions in prompts, with a mean absolute AUC difference of only 0.0084 (0.84%) across all 10 datasets.
-Performance changes are bidirectional, with 5 datasets showing marginal improvements (range: +0.002 to +0.038) and 5 showing marginal declines (range: -0.003 to -0.016) when feature descriptions are included.
-This 50/50 split and small standard deviation (0.0135) indicate no systematic benefit or detriment to either prompt format, suggesting the LLM extracts equivalent semantic information regardless of whether features are referenced by name-description pairs or names alone.
-The method’s robustness to this formatting choice demonstrates that its effectiveness stems from semantic understanding rather than sensitivity to linguistic structure.
-Dataset w/o Desc w/ Desc $\Delta$
-Bank 0.692 0.673 -0.019
-Blood 0.652 0.650 -0.002
-Credit Score 0.756 0.756 +0.000
-Diabetes 0.836 0.836 +0.000
-Heart 0.886 0.893 +0.007
-Income 0.845 0.853 +0.008
-Liver 0.668 0.663 -0.005
-Occupancy 0.989 0.986 -0.002
-Jungle 0.778 0.777 -0.000
-Pima 0.837 0.835 -0.002
-Mean $|\Delta|$ 0.0050
-Table 9: Feature Description Ablation Study: Impact of including feature descriptions in prompts across all datasets. Comparing name-only prompts vs. name+description format.
+We test three variants: (1) concise, (2) verbose, and (3) clinical feature descriptions.
+Examples of these variants are provided in Table 5.
+Results
+Table 6 shows the effect of different grammatical styles for describing features.
+We find that Feature description style shows modest and inconsistent effects.
+Across four datasets, we observe an average performance range of 5.7 AUC points between the best and worst description variants.
+Variant 3 (clinical/formal) achieves the highest mean performance (0.707), outperforming concise (0.679) and verbose (0.664) styles.
+However, the effect is dataset-dependent: Adult Income shows substantial sensitivity (13.1 point range), with Variant 2 performing poorly (0.636 vs.
+0.753/0.767 for other variants), while Stroke exhibits minimal sensitivity (1.9 point range).
+For medical datasets (Stroke, Heart Failure), clinical terminology appears marginally beneficial, suggesting domain-appropriate phrasing may improve prior extraction.
+The inconsistency across datasets indicates that while LLMs are not entirely robust to paraphrasing, the sensitivity is unpredictable and relatively small.
+Variant Example Prompt
+Variant 1 Consider a person who is older. Is this person at risk of stroke? Answer:
+Consider a person who has hypertension. Is this person at risk of stroke? Answer:
+Consider a person who has fewer months since last donation. Does this person have blood donation likelihood? Answer:
+Variant 2 Consider a person who is of older age. Is this person at risk of stroke? Answer:
+Consider a person who has a diagnosis of hypertension. Is this person at risk of stroke? Answer:
+Consider a person who has donated blood more recently. Does this person have blood donation likelihood? Answer:
+Variant 3 Consider a person who exhibits advanced age. Is this person at risk of stroke? Answer:
+Consider a person who presents with hypertension. Is this person at risk of stroke? Answer:
+Consider a person who demonstrates recent donation activity. Does this person have blood donation likelihood? Answer:
+Table 5: Example prompts for feature description variants. All variants use the same template structure but differ in how features are described.
+A.3.2 Contextual preambles
+We test whether providing domain-knowledge context before feature-effect questions improves prior extraction.
+For each feature, we prepend an explanatory preamble describing the causal mechanism (e.g., “Stroke risk increases significantly with age due to arterial stiffening and accumulated vascular damage.
+Based on this medical knowledge,”) before asking the standard question.
+Table 7 shows example prompts with and without contextual preambles.
+Table 6 shows the effect of adding domain-knowledge context before feature-effect questions.
+Feature Description Variants Contextual Preambles
+Dataset Variant 1 Variant 2 Variant 3 Range $\checkmark$ $\times$ $\Delta$
+(Concise) (Verbose) (Clinical) (Baseline)
+Stroke 0.685 0.703 0.704 0.019 0.694 0.707 +0.014
+Blood Donation 0.567 0.582 0.593 0.026 0.583 0.530 -0.053
+Heart Failure 0.713 0.736 0.764 0.051 0.752 0.638 -0.114
+Adult Income 0.753 0.636 0.767 0.131 0.745 0.771 +0.026
+Mean 0.679 0.664 0.707 0.057 0.693 0.662 -0.032
+Table 6: Ablation study results. Left: Effect of feature description style. Right: Effect of contextual preambles.
+Condition Example Prompt
+No Preamble (Baseline) Consider a person who is older. Is this person at risk of stroke? Answer:
+Consider a person who has donated blood more recently. Does this person have blood donation likelihood? Answer:
+Consider a person who has higher ejection fraction. Does this person have poor outcome? Answer:
+With Preamble Stroke risk increases significantly with age due to arterial stiffening and accumulated vascular damage.
+Based on this medical knowledge, consider a person who is older.
+Is this person at risk of stroke?
+Answer:
+People who have donated blood recently are typically more engaged and committed donors.
+Based on this behavioral pattern, consider a person who has donated blood more recently.
+Does this person have blood donation likelihood?
+Answer:
+Higher ejection fraction indicates better cardiac pump function and improved prognosis.
+Based on this cardiac physiology, consider a person who has higher ejection fraction.
+Does this person have poor outcome?
+Answer:
+Table 7: Example prompts comparing baseline (no preamble) with contextual preamble condition.
+Results
+Contextual preambles degrade performance on average.
+Adding domain-knowledge explanations before feature-effect questions reduces mean AUC by 3.2 points (0.693 → 0.662).
+The effect is highly inconsistent: Adult Income (+2.6 points) and Stroke (+1.4 points) benefit modestly from preambles, while Heart Failure (-11.4 points) and Blood Donation (-5.3 points) suffer substantial degradation.
+This suggests that explicit reasoning chains interfere with the LLM’s implicit token probability distributions rather than enhancing them.
+Preambles may bias the model toward generating coherent text completions (which favor the preamble’s stated relationship) rather than revealing the underlying probability mass over answer tokens.
+For prior extraction via logit probabilities, minimal prompting appears superior to verbose contextual framing.
 A.4 Datasets and OOD splits
-A.4.1 Datasets
-We evaluate our prior elicitation method on ten binary, and multi-class classification datasets:
+A.4.1 Public datasets
+Adult Income Becker and Kohavi (1996): Predicts whether an individual’s annual income exceeds \$50K using demographic and educational features such as age, work class, education, marital status, occupation, race, sex, and hours worked per week.
+The binary classification task is based on 48,842 records from the U.S.
+Census Bureau Adult dataset.
+Cervical Cancer Fernandes et al.
+(2017): Predicts the risk of cervical cancer using demographic factors, sexual health history, and medical test results.
+Features include age, number of sexual partners, age at first intercourse, smoking habits, hormonal contraceptive use, and HPV test results.
+The dataset contains 858 patient records collected from clinics in Colombia.
+Chronic Kidney Disease Rubini et al.
+(2015): Classifies the presence of chronic kidney disease using clinical and laboratory measurements, including blood pressure, glucose level, blood urea, serum creatinine, hemoglobin, and albumin.
+The binary classification task is based on 400 patient records from multiple hospitals.
+Bank Marketing Moro et al.
+(2014): Predicts whether a client will subscribe to a term deposit based on demographic attributes and call-related features such as age, job type, marital status, education, balance, contact communication type, and duration of the last call.
+The dataset contains 45,211 records collected from Portuguese banking campaigns.
+Blood Transfusion Service Center Yeh (2008): Predicts whether a donor will donate blood again using historical donation behavior, including recency, frequency, monetary amount of previous donations, and time since first donation.
+The dataset consists of 748 donor records from a blood transfusion center in Taiwan.
+Stroke Fedesoriano (2020): Predicts the risk of stroke using demographic and health-related features such as age, hypertension, heart disease, glucose level, BMI, smoking status, and gender.
+The binary classification task is based on 5110 patient records collected from hospitals in India.
+Heart Failure UCI Machine Learning Repository (2020): Predicts mortality due to heart failure using clinical measurements including age, ejection fraction, serum creatinine, serum sodium, platelets, and smoking status.
+The dataset contains 299 patient records collected from hospitals in Portugal.
+Diabetes Prediction Mustafatz (2023): Classifies diabetes risk using blood glucose levels, BMI, and other demographic and health-related features.
+The dataset contains 520 patient records and is used to predict the onset of diabetes in individuals.
+Give Me Some Credit Kaggle (2011): Predicts the probability of financial distress within two years using demographic and credit-related features such as age, monthly income, number of past delinquencies, number of open credit lines, and debt-to-income ratio.
+The dataset consists of 150,000 records from U.S.
+financial institutions.
+Airfoil Self-Noise Brooks et al.
+(1989): Predicts the noise levels generated by an airfoil using measurements collected by NASA.
+Features include frequency, angle of attack, chord length, free-stream velocity, and suction side displacement thickness.
+The regression task is based on 1,503 measurements of airfoil noise under different operating conditions.
+Bike Sharing Fanaee-T (2013): Predicts hourly bicycle rental demand based on temporal and weather-related features, including hour of the day, day of the week, temperature, humidity, wind speed, and season.
+The dataset consists of 17,379 hourly records collected from a bike-sharing system in Washington, D.C.
+Combined Cycle Power Plant Tfekci and Kaya (2014): Predicts net hourly electrical energy output using ambient variables such as temperature, pressure, humidity, and exhaust vacuum.
+The regression task contains 9,568 records collected from a combined cycle power plant over six years.
+Concrete Compressive Strength Yeh (1998): Predicts the compressive strength of concrete using mixture components such as cement, blast furnace slag, fly ash, water, superplasticizer, coarse aggregate, and fine aggregate, as well as curing age.
+The dataset contains 1,030 records of concrete mixtures tested under controlled conditions.
+Communities and Crime Redmond (2002): Predicts per-capita violent crime rates in U.S.
+communities using demographic, socioeconomic, and law enforcement features, including population, income, employment, education, and police presence.
+The dataset contains 1,994 community records compiled from U.S.
+government sources.
+A.4.2 Private datasets
+Head and Neck Cancer Recurrence Dataset
+In collaboration with Henry Ford Health, we compiled a binary classification dataset for Head and Neck Cancer Recurrence.
+This dataset includes comprehensive features about patient demographics, tumor characteristics, treatment regimens, and clinical staging, and predicts the likelihood of cancer recurrence following radiotherapy.
+The dataset encompasses 508 patients with head and neck cancers, capturing clinical variables including comorbidity indices (Charlson score), smoking history (pack-years), HPV status, primary tumor site localization, and detailed radiation therapy parameters such as treatment duration, prescription dose, and the critical time interval between biopsy and treatment initiation.
+This dataset presents a significant class imbalance (390 non-recurrence vs.
+118 recurrence cases), reflecting real-world clinical distributions.
+The dataset is particularly valuable for evaluating model performance under distributional shift, as we created an out-of-distribution test set focusing on elderly patients ($age\geq 75$), representing a clinically meaningful subpopulation that is not present in any LLM pre-training corpora.
+Moreover, this dataset highlights why our approach is necessary.
+Due to its limited sample size, some patterns learned from the data are not medically plausible.
+Table 8 shows several noticeable relationships between features and the target variable.
+For instance, the data suggests that marijuana use is associated with a lower risk of Head and Neck Cancer recurrence, which contradicts established medical evidence Kim et al.
+(2020).
+This discrepancy is likely due to sampling bias or confounding factors in the small dataset.
+LoID priors address this issue by incorporating domain knowledge into the model.
+In this case, they can enforce a prior that reflects a more realistic relationship between marijuana use and cancer recurrence.
+As a result, the model is guided away from spurious correlations, making it more reliable and better suited for generalization to real-world populations.
+Recurrence Yes No
+Sample Size 118 390
+CONTINUOUS FEATURES (Mean [Standard Deviation])
+Pack Years 41.3 (38.1) 27.3 (28.7)
+Biopsy-RT Duration 44.6 (28.2) 46.2 (47.5)
+Average Income 82348.7 (32140.4) 90597.7 (36881.5)
+CATEGORICAL FEATURES (Mode [%])
+Clinical N 2a (50.0) 0 (29.0)
+Clinical T 2 (53.7) 2 (57.6)
+Radiation Type IMRT Standard (50.4) IMRT Vmat (52.2)
+Alcohol Frequent (41.3) No Alcohol Use (45.8)
+Smoking History Former Use (45.4) Never Smoker (41.5)
+Grade of Differentiation Moderate (68.7) Moderate (79.0)
+Neck RT Bilateral (85.1) Bilateral (91.7)
+Tobacco Yes (54.5) No (63.2)
+Marijuana No (87.6) No (73.5)
+First Primary Site = Larynx-supraglottis No (73.6) No (82.0)
+First Primary Site = Larynx-glottis No (82.6) No (92.0)
+First Primary Site = Oropharynx-base of tongue No (78.5) No (69.5)
+First Primary Site = Oropharynx-tonsil No (79.3) No (70.5)
+Concurrent Chemotherapy No (90.1) No (95.5)
+Table 8: Notable feature–target relationships in the Head and Neck Cancer Recurrence dataset.
+While some associations, such as those involving tobacco use, pack years, and alcohol consumption, align with clinical expectations, others are misleading.
+In particular, a larger proportion of marijuana users appear in the non-recurrence group compared to the recurrence group, which may incorrectly suggest a protective effect.
+A standard logistic regression model may learn this spurious correlation, despite it lacking medical validity.
+A.5 Out-of-Distribution splits
+Our goal in this research is to evaluate how well a model trained on a small, specific population generalizes to a larger population.
+To simulate this, we create conditions where only a subset of the full population is available for training, but the aim remains to extract meaningful patterns from the data.
+Specifically, we first generate train and test sets representing the entire population using a standard 80-20 shuffle split.
+From the training set, we then select a subset that excludes certain samples entirely, creating a subpopulation that is not fully representative of the overall population.
+This setup reflects real-world scenarios, such as Head and Neck Cancer Recurrence, where the limited number of training samples can hinder pattern discovery.
+Table 9 summarizes the split criteria for all datasets.
+Binary Classification:
+  • Adult (Income Prediction): Training set excludes individuals age <28<28.
+  • Chronic Kidney Disease: Training set excludes patients with blood urea <61.75<61.75 mg/dL.
+  • Cervical Cancer: Training set excludes women whose age at first sexual intercourse $\leq 15$.
+  • Blood Donation: Training set excludes donors with recency >14>14 months (i.e., recent donors are excluded).
+  • Stroke: Training set includes only patients age $\leq 32$.
+  • Heart Failure: Training set includes only patients with ejection fraction $\geq 30\%$.
+  • Diabetes Prediction: Training set includes patients with blood glucose $\leq 126$ mg/dL; testing set includes hyperglycemic patients with glucose $\geq 126$ mg/dL.
+  • Bank Marketing: Training set excludes calls with duration >127>127 seconds.
+  • Head & Neck Cancer (HNC): Training set includes patients age $\leq 75$.
+  • Good Credit Score: Training set excludes individuals age >31>31.
+Regression:
+  • Airfoil Noise: Training set includes low-frequency conditions (frequency <1600<1600 Hz).
+  • Bike Sharing: Training set includes early hours (hour $\leq 6$).
+  • Combined Cycle Power: Training set includes moderate temperatures (AT $\leq 13.51^{\circ}$C).
+  • Communities & Crime: Training set includes racially diverse areas (racePctWhite $\leq 0.98$).
+  • Concrete Strength: Training set includes high-cement mixes (>350>350 kg/m³), where mixture interactions change.
+Dataset Type Split Feature OOD Test Condition $P_{\text{IID}}$ $P_{\text{OOD}}$
+Binary Classification
+Adult Binary Age Age << 28 0.809 0.414
+Cervical Cancer Binary First sexual intercourse Age << 15 0.870 0.440
+Chronic Kidney Disease Binary Blood urea BU << 61.75 mg/dL 0.993 0.822
+Diabetes Prediction Binary Blood glucose Glucose >> 126 mg/dL 0.963 0.883
+Stroke Binary Age Age >> 32 0.842 0.608
+Heart Failure Binary Ejection fraction EF << 30% 0.849 0.746
+Bank Marketing Binary Call duration Duration >> 127 sec 0.878 0.780
+Blood Donation Binary Recency Recency >> 14 months 0.748 0.572
+Head & Neck Cancer Binary Age Age >> 75 0.998 0.542
+Good Credit Score Binary Age Age >> 31 0.872 0.800
+Regression
+Concrete Strength Regression Cement Cement >> 350 kg/m³ 98.1 469.2
+Communities Crime Regression Race (White %) White >> 98% 0.007 0.171
+Airfoil Noise Regression Frequency Freq << 1600 Hz 23.7 77.2
+Combined Cycle Power Regression Temperature Temp >> 13.51°C 21.2 962.9
+Bike Sharing Regression Hour Hour >> 6 19030 77858
+Table 9: Out-of-distribution split specifications for all datasets. Train/test splits are created by partitioning on a single feature to induce distributional shift with non-overlapping support.
+A.5.1 Rationale
+These splits target realistic distribution shifts where:
   1. 1.
-    Bank Marketing Moro et al.
-(2014): Predicts whether a client will subscribe to a term deposit based on direct marketing campaign data from a Portuguese banking institution.
-Features include client demographics (age, job, marital status, education), campaign information (contact type, number of contacts, previous campaign outcome), and economic indicators (employment variation rate, consumer price index).
-The binary classification task identifies clients likely to subscribe to term deposits.
+    Subpopulation exclusion is plausible: Training data may naturally lack certain demographics (young adults in income surveys), extreme operating conditions (high-frequency turbine tests), or rare clinical presentations (severe heart failure cases).
   2. 2.
-    Blood Transfusion Yeh (2008b): Predicts whether a blood donor will donate within a given time window using data from the Blood Transfusion Service Center in Hsin-Chu City, Taiwan.
-Features are based on the RFMTC marketing model: Recency (months since last donation), Frequency (total number of donations), Monetary (total blood donated in c.c.), and Time (months since first donation).
-The binary classification task identifies donors likely to donate in March 2007.
+    Feature relationships may differ: Age-income correlations differ for young adults vs. middle-aged workers; thermodynamic equations behave differently at temperature extremes.
   3. 3.
-    Give Me Some Credit Kaggle (2011): Predicts the probability of financial distress in the next two years using credit and demographic data from Kaggle.
-Features include checking account status, credit duration, credit history, credit amount, savings account balance, employment status, and age.
-The binary classification task evaluates financial risk for 1,000 borrowers.
-  4. 4.
-    Diabetes Islam et al.
-(2020): Predicts the likelihood of diabetes based on self-reported symptoms and demographic information.
-Features include age, gender, polyuria, polydipsia, sudden weight loss, weakness, polyphagia, genital thrush, visual blurring, itching, irritability, delayed healing, partial paresis, muscle stiffness, alopecia, and obesity.
-The dataset contains 520 instances collected via questionnaires and is formulated as a binary classification task.
-  5. 5.
-    Heart Disease Janosi et al.
-(1988): Predicts heart disease using clinical and demographic attributes such as age, sex, resting blood pressure, cholesterol level, fasting blood sugar, maximum heart rate, exercise-induced angina, ST depression (oldpeak), chest pain type, resting ECG results, and ST slope.
-The binary classification task identifies individuals with high likelihood of heart disease.
-  6. 6.
-    Adult Income Becker and Kohavi (1996a): Predicts whether an individual’s annual income exceeds \$50K based on 1994 U.S.
-Census data.
-Features include age, education level, marital status, occupation, relationship status, workclass, sex, native country, capital gain/loss, and hours worked per week.
-The binary classification task uses demographic and employment data from 48,842 records.
-  7. 7.
-    Chess Endgame Shapiro (1983): Predicts game outcomes in King-Rook vs.
-King-Pawn chess endgame positions.
-Features encode piece positions using strength rankings, file coordinates, and rank coordinates for white and black pieces.
-The binary classification task determines winning positions in this chess endgame with 44,820 board configurations.
-  8. 8.
-    Indian Liver Patient Ramana and Venkateswarlu (2012): Predicts whether a patient has liver disease using clinical measurements collected from northeast Andhra Pradesh, India.
-Features include age, gender, total bilirubin, direct bilirubin, alkaline phosphatase, alanine aminotransferase (SGPT), aspartate aminotransferase (SGOT), total proteins, albumin, and albumin/globulin ratio.
-The binary classification task distinguishes liver patients from healthy individuals using data from 583 records (416 liver patients, 167 non-liver patients).
-  9. 9.
-    Occupancy Detection Candanedo (2016a): Predicts room occupancy status from environmental sensor measurements in an office setting.
-Features include temperature (Celsius), relative humidity (%), light (Lux), CO2 concentration (ppm), humidity ratio (kg water-vapor/kg-air), and timestamp information.
-The binary classification task detects occupancy using measurements recorded at one-minute intervals.
-  10. 10.
-    Pima Indians Diabetes Smith et al.
-(1988): Predicts the onset of diabetes within five years in females of Pima Indian heritage over age 21.
-Features include number of pregnancies, glucose level, blood pressure, skin thickness, insulin level, BMI, diabetes pedigree function, and age.
-The binary classification task uses medical measurements from 768 women from a population near Phoenix, Arizona with high diabetes incidence rates.
-Dataset Method Mean Mag. Std (Uncertainty) Std Ratio Performance
-[…]
-AutoElicit 0.051 0.058 $1\times$
-[…]
-AutoElicit 0.015 0.124 $1\times$
-[…]
-AutoElicit 0.017 1.379 $1\times$
-[…]
-AutoElicit 0.008 0.047 $1\times$
-[…]
-AutoElicit 0.015 0.087 $1\times$
-[…]
-AutoElicit 0.070 0.044 $1\times$
-[…]
-AutoElicit 0.085 0.059 $1\times$
-[…]
-AutoElicit 0.007 0.122 $1\times$
-[…]
-AutoElicit 0.025 0.114 $1\times$
-[…]
-AutoElicit 0.041 0.024 $1\times$
-Table 10: Comparison of LoID and AutoElicit across datasets in terms of magnitude, uncertainty, and performance.
-A.4.2 OOD Splits
-We construct out-of-distribution (OOD) train splits to simulate a training set that does not represent the entire population.
-For each dataset, we consider numerical features with at least two unique values and apply five shift strategies to induce distributional gaps between training and test data:
-  1. 1.
-    extreme_10, training on the bottom 10%
-  2. 2.
-    extreme_5_95, training on the bottom 5%
-  3. 3.
-    moderate_20_80, training on the bottom 20%;
-  4. 4.
-    tail_0_50, training on the 0–25% range;
-  5. 5.
-    tail_50_100, training on the 50–75% range.
-For each split, we designate a different feature as the shift variable (cycling through available features), compute the corresponding quantile thresholds, and construct train–test masks while enforcing a minimum of 50 samples per split.
-Models are trained on the OOD training set and evaluated on the entire dataset to quantify performance degradation under distribution shift.
-Splits with insufficient class diversity (fewer than two classes in either the training or test set) are discarded.
-This procedure simulates realistic deployment settings in which models must generalize to regions of the feature space not observed during training.
-For Heart Disease, we use the cholesterol split (moderate_20_80), training on the 20–80% quantile range and testing on extreme values, resulting in an AUC drop from 0.933 to 0.870 ($\Delta=0.063$).
-Liver Disease uses the alkaline phosphatase split (tail_0_50), training on the bottom 50% and testing on the top 50%, with degradation from 0.768 to 0.698 ($\Delta=0.070$).
-Blood Donation employs the recency split (extreme_10_90), training on donors in the bottom 10% and testing on the top 10%, experiencing the largest degradation from 0.755 to 0.579 ($\Delta=0.176$).
-Occupancy Detection uses the light intensity split (moderate_20_80), although this split exhibits insufficient class diversity for full evaluation.
-Diabetes and Pima both use the pregnancies split (extreme_10_90), training on the bottom 10% and testing on the top 10%, with degradation from 0.839 to 0.813 ($\Delta=0.027$).
-CreditG partitions on credit amount (extreme_5_95), training on the bottom 5% and testing on the top 5%, yielding a drop from 0.786 to 0.710 ($\Delta=0.076$).
-Income uses the education level split (extreme_5_95), with the steepest decline from 0.908 to 0.781 ($\Delta=0.127$).
-Bank Marketing splits on call duration (extreme_10_90), experiencing severe degradation from 0.888 to 0.508 ($\Delta=0.380$), representing the most extreme covariate shift in our evaluation.
-Finally, Jungle Chess uses the black piece strength split (tail_0_50), training on the 0–50% range and testing on 50–100%, with degradation from 0.810 to 0.712 ($\Delta=0.098$).
-These splits simulate realistic deployment scenarios in which models trained on one subpopulation must generalize to substantially different data distributions.
-[…]
-Bank 0.51 0.61 0.72 0.61 0.89 +30.0
-Blood 0.58 0.63 0.67 0.62 0.75 +31.3
-Credit Score 0.71 0.74 0.76 0.76 0.79 +0.0
-Diabetes 0.81 0.82 0.84 0.81 0.84 +104.5
-Heart 0.87 0.91 0.90 0.92 0.93 -31.8
-Income 0.78 0.81 0.85 0.83 0.91 +14.9
-Liver 0.70 0.68 0.69 0.70 0.77 -9.9
-Occupancy 0.98 0.72 1.00 0.75 1.00 +1155.7
-Jungle 0.71 0.72 0.79 0.75 0.81 +40.0
-Pima 0.82 0.81 0.84 0.81 0.84 +98.1
-Average 0.74 0.75 0.80 0.76 0.86 +143.3
-Table 11: Comparison of LoID priors and priors with 0 means and stds in the Goldilock zone on OOD datasets.
-$\Delta$ Gap Closed shows the percentage point difference in gap closed between LoID and Prior = $\mathcal{N}=(0,0.45)$.
-OOD-LR is logistic regression trained on OOD data (lower bound), Cap is logistic regression trained on full IID data (upper bound).
-A.5 LoID vs Existing Baselines
-A.5.1 Prior Performance
-Since LoID and AutoElicit employ similar prior elicitation strategies and exhibit more comparable performance than other baselines, Table 10 focuses on analyzing where and why each method performs better.
-The underlying intuition behind these differences is discussed in detail in Section 5.1.
-A.5.2 Computational Efficiency
+    Prior knowledge should help: LLMs may have generalizable knowledge (e.g., "older age increases stroke risk") that transfers to the OOD regime, even if training data provides no direct evidence.
+A.6 LoID vs. correlation analysis
+In addition to the baselines mentioned in Table 1, we test whether simple correlation analysis might capture similar relationships to LLM priors.
+To address this, we test a correlation-based baseline that uses feature-target Pearson correlations from the training data as prior means (with fixed $\sigma=1$).
+Table 10 compares LoID against correlation-based priors.
+As expected, correlations learned from OOD train sets can mislead the model.
+In the Bank Marketing dataset where LoID failed because it is dataset-specific and in Cumbine Cycle Power where $\mathcal{N}(0,1)$ could also regain a lot of the performance showing that the train set was actually kinda representing the test set and so correlation studies also got a good AUC.
+These results further confirm our previous findings.
+Binary Classification (AUC) Regression (MSE)
+Dataset LoID Corr. $\Delta$ Dataset LoID Corr. $\Delta$
+Adult 0.725 0.529 -0.196 Airfoil Noise 60.5 51.4 -9.1
+Bank Marketing 0.652 0.770 +0.118 Bike Sharing 78924 69969 -8955
+Blood Donation 0.663 0.574 -0.089 Combined Cycle Power 961.9 24.1 -937.8
+Cervical Cancer 0.681 0.450 -0.231 Communities & Crime 0.07 0.18 +0.11
+Chronic Kidney Disease 0.901 0.575 -0.326 Concrete Strength 121.0 2292.2 +2171.2
+Diabetes Prediction 0.906 0.873 -0.033
+Heart Failure 0.753 0.742 -0.011
+HNC 0.696 0.566 -0.130
+Stroke 0.692 0.554 -0.138
+Good Credit Score 0.760 0.376 -0.382
+Mean 0.741 0.626 -0.142 Mean 16013 14467 -1546
+Table 10: LoID (LLM priors) vs.
+correlation-based priors across 13 datasets (8 classification, 5 regression).
+For regression, lower MSE is better; negative $\Delta$ indicates correlation outperforms LoID.
+A.7 Computational efficiency
 We compare the computational complexity and runtime of LoID with AutoElicit and LLMProcesses.
@@
 $T_{\text{AE}}=\mathcal{O}(d\times N_{\text{rounds}}\times N_{\text{queries}}\times T_{\text{LLM}}),$
-where $N_{\text{rounds}}\in[3,5]$ and $N_{\text{queries}}\in[5,10]$. This results in more than 28 LLM queries per feature on average.
+where $N_{\text{rounds}}=10$ and $N_{\text{queries}}=10$. This results in more than 28 LLM queries per feature on average.
 […]
@@
 Unlike LoID and AutoElicit, which perform prior elicitation once per dataset, LLMProcesses scales linearly with the test set size and incurs substantially longer runtimes (15–20 hours for datasets with 100–400 test instances, compared to minutes for prior elicitation methods).
-A.6 Prior Impact on Model Performance
-A.6.1 LoID Priors vs. $\mathcal{N}$(0, 0.45) vs. $\mathcal{N}$(0, 1)
-To isolate the contribution of the prior means inferred by LoID, we evaluate a baseline using normal priors with zero mean and standard deviation $0.45$ (the Goldilock setting).
-This comparison quantifies how much LoID’s learned prior means contribute to downstream AUC.
-As shown in Table 11, incorporating LoID-derived means yields a substantial performance gain, increasing the gap closed by approximately 154% on average.
-This difference is most prominent in datasets where the uninformed priors catastrophically fail: Occupancy (+1156%), Diabetes (+105%), and Pima (+98%).
-In these cases, the $\mathcal{N}(0,0.45)$ prior provides insufficient directional guidance, causing the Bayesian posterior to converge toward coefficients near zero and perform worse than the OOD baseline.
-Conversely, LoID’s informative means—despite systematic positive bias—capture feature importance magnitudes that guide the model toward the correct solution space.
-Even on datasets where LoID underperforms (Heart: -32%, Liver: -10%), the performance gap remains small compared to the massive gains elsewhere.
-This demonstrates that prior mean quality matters substantially more than perfect directional correctness when paired with appropriately calibrated uncertainty ($\sigma\approx 0.4$$0.5$).
-Finally, we include an additional $\mathcal{N}(0,1)$ baseline to illustrate the improvement achieved over a more weakly informative prior that is commonly used with Bayesian inference.
-A.6.2 $\mathcal{N}$(0, 1) vs. $\mathcal{U}$(-1, 1)
-To understand the impact of prior distribution choice, we compare two configurations: (1) uninformative Normal priors $\mathcal{N}(0,1)$, and (2) uninformative Uniform priors $\mathcal{U}(-1,1)$.
-Table 12 shows the catastrophic failure of Uniform priors (average AUC 0.42, with Occupancy at 0.02) and reveals that while Uniform priors ostensibly provide “maximum uninformativeness” by assigning equal probability across a range, they fundamentally fail as regularizers for regression coefficients because:
-  1. 1.
-    Lack of soft regularization: Uniform distributions do not encode the reasonable prior belief that smaller coefficient magnitudes are more plausible than larger ones.
-This absence of gradient-based regularization toward zero can lead to poor MCMC sampling behavior and overfitting.
-  2. 2.
-    Incompatibility with MCMC: The flat density of Uniform priors provides no gradient information to guide sampling algorithms, resulting in inefficient exploration and convergence issues, particularly visible in the Occupancy dataset collapse.
-  3. 3.
-    Prior-likelihood mismatch: For logistic regression, coefficients naturally follow a scale where values near zero are common and extreme values are rare.
-Uniform priors violate this structure, treating $\beta=0.1$ and $\beta=4.9$ as equally plausible a prior.
-Dataset $\mathcal{N}(0,1)$ $\mathcal{U}(-1,1)$
-Bank 0.61 0.35
-Blood 0.63 0.59
-Credit Score 0.74 0.58
-Diabetes 0.82 0.32
-Heart 0.91 0.84
-Income 0.81 0.30
-Liver 0.68 0.70
-Occupancy 0.72 0.02
-Jungle 0.72 0.58
-Pima 0.82 0.32
-Average 0.75 0.46
-Table 12: Comparison of LoID priors against uninformative Normal and Uniform priors on OOD datasets.
-A.7 Computational Resources
+A.8 Resources and reproducibility
+Code availability
+To ensure reproducibility, we will release our full codebase, including detailed instructions for running the experiments and reproducing all reported results.
+We also provide the the posterior files generated by our experiments, as well as the scripts used to generate all figures and tables in the paper.
+Data availability
+All datasets used in this study except Head and Neck Cancer Recurrence are publicly available through open-access repositories.
+Further details and references are provided in Section A.4.
+All preprocessing steps required to obtain the final train/test splits used in our experiments, along with the preprocessed data, are available in our code release.
+The Head and Neck Cancer Recurrence dataset will be released upon publication.
+Results variability
+To assess the stability and determinism of our approach, all experiments were repeated five times.
+The resulting 95% confidence intervals for AUC were effectively zero (0.000), indicating that our method yields highly consistent results across runs.
+Computational resources
   1. 1.